### v2

# Arithmetical Complexity and Absoluteness of Rigidity Phenomena for Ulam Sequences

Abstract
We analyse the logical complexity and absoluteness of natural statements about Ulam sequences, with particular emphasis on the rigidity phenomena introduced by Hinman, Kuca, Schlesinger and Sheydvasser for the family $U(1,n)$. For each pair of coprime integers $a<b$ we view the associated Ulam sequence $U(a,b)$ as a recursive subset of $\mathbb{N}$ and consider expansions of the form $(\mathbb{N},+,\mathrm{U}_{a,b})$.
Our first main result is a uniform coding of Ulam sequences and of the interval-with-periodic-mask patterns appearing in rigidity conjectures into first-order arithmetic. Using this, we show that strong rigidity, regularity (eventual periodicity of gaps), and density statements for $U(a,b)$ are all arithmetical and lie at low levels of the arithmetical hierarchy (specifically $\Sigma^{0}_{2}$ or $\Pi^{0}_{3}$).
As a consequence, these statements are absolute between transitive models of $\mathrm{ZFC}$ with the same natural numbers: their truth value cannot be changed by forcing, and is independent of the Continuum Hypothesis and large cardinal axioms. We also study the expansions $(\mathbb{N},+,\mathrm{U}_{a,b})$ model-theoretically, showing that combinatorial rigidity implies tameness (NIP, dp-minimality, non-interpretability of multiplication).
keywords:
Ulam sequences , rigidity , gap regularity , asymptotic density , arithmetical hierarchy , absoluteness , Presburger arithmetic , NIP , dp-minimality
2020 MSC:
11B83 , 03F25 , 03C64 , 11B05 , 03E35
1 Introduction
1.1 Ulam sequences and rigidity
In his 1964 paper [Ulam1964], Ulam introduced what is now called the Ulam sequence $U(1,2)$ as a simple experiment in additive number theory: start with $1$ and $2$, and then recursively add the smallest integer that can be written as a sum of two distinct earlier terms in a unique way. Thus
$1,2,3,4,6,8,11,13,16,18,26,28,\dots$
is obtained by admitting $3=1+2$, then $4=1+3$, while skipping $5$, which can be written both as $1+4$ and $2+3$. More generally, for integers $a<b$ one defines the Ulam sequence $U(a,b)$ by starting from $\{a,b\}$ and enforcing the same “unique representation as a sum of two distinct earlier terms” rule.
Despite the simplicity of this definition, Ulam sequences have proved remarkably resistant to analysis. Even for the classical sequence $U(1,2)$, fundamental basic questions remain open: for example, it is not known whether $U(1,2)$ has a natural density, or whether its gap sequence is eventually periodic. A series of works by Finch, Schmerl–Spiegel and Cassaigne–Finch [Finch1992ZeroAdditive, SchmerlSpiegel1994, CassaigneFinch1995] initiated a systematic study of more general “$1$-additive” sequences, including $U(a,b)$, and identified many examples in which the gap sequence eventually becomes periodic. In particular, they showed that for a number of pairs $(a,b)$ the sequence $U(a,b)$ has only finitely many even terms and is therefore regular in this sense.
A second strand of work has highlighted striking global patterns in Ulam sequences. Steinerberger [Steinerberger2017HiddenSignal] observed that for a certain explicit real number $\alpha$, the sequence of fractional parts $\{u_{n}\alpha\}$, where $(u_{n})$ enumerates $U(1,2)$, exhibits a highly non-uniform limiting distribution concentrated in the middle third of $[0,1]$. This “hidden signal” is especially surprising in light of Weyl’s equidistribution theorem, which asserts that for any sequence of distinct integers $(a_{n})$ and Lebesgue-almost every $\alpha$, the sequence $\{a_{n}\alpha\}$ is equidistributed.
Building on these observations, Hinman, Kuca, Schlesinger and Sheydvasser [HinmanKucaSchlesingerSheydvasser2019JNT, HinmanKucaSchlesingerSheydvasser2019Involve] developed a series of rigidity results for Ulam sequences. Roughly speaking, they showed that as one varies the second term $n$ in $U(1,n)$, large initial segments of the resulting sequences can be described by a finite amount of data that does not depend on $n$ itself, but only on its residue class modulo some fixed $L$. More concretely, for the family $U(1,n)$ they produce decompositions of $U(1,n)\cap[1,c_{j}n+d_{j}]$ (where $c_{j}\to\infty$), valid for all sufficiently large $n$ in a fixed congruence class, as finite unions of intervals with endpoints depending linearly on $n$.
Remark 1.1 (Status of rigidity for $U(a,b)$ with $a>1$).
The rigidity results of [HinmanKucaSchlesingerSheydvasser2019JNT] are established for the family $U(1,n)$. For sequences $U(a,b)$ with $a>1$, computational evidence collected since the original rigidity paper suggests that strong rigidity in the same form does not hold in general, although the correct conjectural description remains unclear (A. Sheydvasser, personal communication, 2025). We therefore focus primarily on the $U(1,n)$ family when discussing proven rigidity phenomena, while formulating our arithmetical complexity results for general coprime pairs $(a,b)$. The complexity classifications themselves (Theorems 4.35.8) hold regardless of whether the corresponding rigidity or regularity statements are true.
These results support the view that Ulam sequences, while seemingly irregular at small scales, exhibit a surprising degree of large-scale order—at least in the $U(1,n)$ family.
1.2 Main results and significance
The central contribution of this paper is to determine the logical complexity of statements about Ulam sequences and to draw consequences from this classification. Our main theorem, stated informally, is:
Main Theorem (Informal). The strong rigidity conjecture, gap regularity, and density statements for Ulam sequences $U(a,b)$ are all arithmetical statements of low complexity (specifically $\Sigma^{0}_{2}$ or $\Pi^{0}_{3}$). As a consequence:
  1. 1.
    Adding or removing set-theoretic axioms such as the Axiom of Choice, the Continuum Hypothesis, or large cardinal axioms does not change their truth values.
  2. 2.
    No set-theoretic forcing extension of $\mathrm{ZFC}$ can make one of these statements true while another such extension (with the same natural numbers) makes it false.
  3. 3.
    If $U(a,b)$ satisfies gap regularity, then the expansion $(\mathbb{N},+,\mathrm{U}_{a,b})$ is model-theoretically tame: NIP, dp-minimal, and does not interpret multiplication.
Why this matters: for number theorists
The rigidity conjectures for Ulam sequences are purely combinatorial questions about sequences of integers. It is natural to ask whether sophisticated set-theoretic machinery might help resolve them—perhaps the answer depends on whether we accept certain axioms, or perhaps forcing techniques could shed light on the problem.
Our results show definitively that this cannot happen. The statements in question are absolute: their truth or falsity is already determined by the structure of the natural numbers, and no set-theoretic axiom or technique can change the answer. This is both reassuring (the answer does not depend on one’s foundational commitments) and sobering (there are no set-theoretic shortcuts to resolving these combinatorial questions).
More specifically, we show that the key conjectures can be expressed using only two or three alternations of quantifiers over natural numbers. This places them at a very low level of logical complexity—comparable to statements like “there exist infinitely many twin primes” or “Goldbach’s conjecture holds.” The classification gives a precise sense in which these problems, while difficult, are “elementary” from a logical standpoint.
Why this matters: for model theorists
Ulam sequences provide a natural family of recursively defined predicates on $\mathbb{N}$ whose model-theoretic behavior is controlled by their combinatorial structure. The expansion $(\mathbb{N},+,\mathrm{U}_{a,b})$ sits between two well-understood extremes: Presburger arithmetic $(\mathbb{N},+)$, which is decidable and tame, and full arithmetic $(\mathbb{N},+,\times)$, which is undecidable and wild.
Under the assumption of combinatorial rigidity, we show that these expansions remain on the tame side: they are definitional expansions of Presburger arithmetic, hence NIP and dp-minimal. Without assuming rigidity, the situation is open and provides concrete test cases for questions about expansions of $(\mathbb{N},+)$ by sparse additive sets. The paper initiates a systematic study of these expansions and formulates precise conjectures about their tameness.
1.3 Logical background for non-logicians
For readers unfamiliar with mathematical logic, we provide informal explanations of the key concepts that appear in our results. Precise definitions are given in Section 2.
The arithmetical hierarchy
The arithmetical hierarchy classifies statements about natural numbers by the pattern of quantifiers (“for all” and “there exists”) needed to express them. The notation $\Sigma^{0}_{n}$ and $\Pi^{0}_{n}$ indicates $n$ alternating blocks of quantifiers:
$\Sigma^{0}_{1}$: $\exists x\,\varphi(x)$ where $\varphi$ is “directly checkable”
$\Pi^{0}_{1}$: $\forall x\,\varphi(x)$ where $\varphi$ is “directly checkable”
$\Sigma^{0}_{2}$: $\exists x\,\forall y\,\varphi(x,y)$
$\Pi^{0}_{2}$: $\forall x\,\exists y\,\varphi(x,y)$
$\Sigma^{0}_{3}$: $\exists x\,\forall y\,\exists z\,\varphi(x,y,z)$
Lower levels correspond to “simpler” statements. For example, “$U(1,2)$ has infinitely many even elements” is $\Pi^{0}_{2}$ (for all $N$, there exists an even $m>N$ in $U(1,2)$), while “the gap sequence of $U(a,b)$ is eventually periodic” is $\Sigma^{0}_{2}$ (there exist $N$ and $p$ such that for all $k\geq N$, the gaps repeat with period $p$).
Absoluteness
A statement is absolute if its truth value does not depend on the ambient set-theoretic universe. Different models of set theory (ZFC) can disagree about many things—the size of the continuum, whether certain large cardinals exist, etc.—but they all agree on arithmetical statements, because they all contain the same natural numbers with the same addition and multiplication.
The practical consequence: if you prove that a conjecture about Ulam sequences is arithmetical (as we do), then you know immediately that the Axiom of Choice, the Continuum Hypothesis, large cardinal axioms, and set-theoretic forcing cannot affect its truth value.
NIP and tameness
In model theory, structures are classified by how “wild” or “tame” their definable sets are. A theory is NIP (“not the independence property”) if its definable families of sets cannot encode arbitrary combinatorial patterns—roughly, there is a uniform bound on how complicated the interactions between definable sets can be.
Presburger arithmetic $(\mathbb{N},+)$ is a paradigm of tameness: NIP, decidable, and with a complete understanding of definable sets (they are “semilinear”). Full arithmetic $(\mathbb{N},+,\times)$ is wild: it can encode arbitrary computation, and its definable sets are extremely complicated.
When we say that $(\mathbb{N},+,\mathrm{U}_{a,b})$ is “tame” under rigidity assumptions, we mean it inherits the good behavior of Presburger arithmetic rather than the pathologies of full arithmetic.
1.4 Technical overview
The starting point of this paper is the observation that all of the objects above—the individual Ulam sequences $U(a,b)$, the families $\{U(a,b)\}$, and the associated rigidity patterns—are inherently arithmetical. For each coprime pair $a<b$, membership in $U(a,b)$ is algorithmically decidable: given $(a,b,m)$, one can simulate the Ulam construction up to $m$ and check whether $m$ is admitted. In particular, each $U(a,b)$ is a recursive subset of $\mathbb{N}$, and there is a uniform $\Delta^{0}_{1}$ predicate $\mathrm{U}(a,b,m)$ expressing $m\in U(a,b)$ in first-order arithmetic. Likewise, the “finite patterns” appearing in rigidity conjectures can be encoded by natural numbers and decoded by primitive recursive functions.
From this perspective, Ulam sequences give rise to a large family of first-order structures
$\mathcal{U}_{a,b}\;=\;(\mathbb{N},+,0,1,\mathrm{U}_{a,b}),$
where $\mathrm{U}_{a,b}(m)$ abbreviates $\mathrm{U}(a,b,m)$. Natural questions about rigidity, regularity and density then become questions about the first-order theory of these expansions and the logical form of particular sentences in the language of arithmetic. For example:
  • the assertion that $U(a,b)$ is eventually a finite union of linearly parameterised intervals with a periodic mask becomes a sentence with a specific pattern of first-order quantifiers over $\mathbb{N}$;
  • the assertion that the gap sequence of $U(a,b)$ is eventually periodic can be written as $\exists N\exists p\forall k\geq N\,(\dots)$, and thus occupies a definite place in the arithmetical hierarchy;
  • statements asserting the existence and value of the asymptotic density of $U(a,b)$ can likewise be analysed in terms of their quantifier complexity.
Once one knows where these statements sit in the arithmetical hierarchy, a number of consequences follow almost for free. In particular, purely arithmetical sentences are absolute between transitive models of $\mathrm{ZFC}$ with the same natural numbers, and their truth value cannot be changed by forcing extensions. Thus, if a strong rigidity statement about Ulam sequences turns out to be, say, $\Sigma^{0}_{2}$ or $\Pi^{0}_{3}$, then it cannot depend on the Continuum Hypothesis, forcing axioms, or large cardinal hypotheses: whatever its truth value is, it is already determined inside first-order arithmetic. This provides a sharp separation between the combinatorial substance of the rigidity phenomena and the ambient set-theoretic background.
Beyond absoluteness, one can also ask about the model-theoretic behaviour of the expansions $\mathcal{U}_{a,b}$. A priori such expansions need not be tame: for a general recursive predicate $P\subseteq\mathbb{N}$, the structure $(\mathbb{N},+,P)$ may interpret arbitrary complicated combinatorics, including multiplication, and may have the independence property (IP) in the sense of Shelah[ShelahCT] (for these notions, see Section 7.1). However, the empirical and rigorous rigidity results for Ulam sequences suggest that the specific predicates $\mathrm{U}_{a,b}$ could be much better behaved. One of the themes of this paper is that strong combinatorial rigidity of $U(a,b)$ should correlate with model-theoretic tameness of $\mathcal{U}_{a,b}$, for instance with the absence of IP or with non-interpretability of $(\mathbb{N},\times)$.
The aim of this work is therefore twofold. First, we give a uniform arithmetisation of Ulam sequences and of the “finite patterns” used in rigidity results, and we identify the precise arithmetical complexity of a range of natural statements about $U(a,b)$: strong rigidity, regularity, finiteness of even terms, and various density assertions. Second, we draw out the resulting absoluteness consequences and outline some initial model-theoretic implications for the structures $\mathcal{U}_{a,b}$. In this way, we place the rigidity phenomena for Ulam sequences within the broader landscape of definability, absoluteness and classification theory in mathematical logic.
1.5 Main results
We now summarise our main contributions more formally. Throughout, we write $\mathrm{U}(a,b,m)$ for the uniform $\Delta^{0}_{1}$ predicate coding membership in $U(a,b)$, as constructed in Theorem 3.2, and $\mathrm{InPattern}(c,a,b,m)$ for the uniform primitive recursive relation expressing that $m$ belongs to the pattern encoded by $c$ for parameters $(a,b)$.
Uniform arithmetisation.
Our first result is a uniform arithmetisation of Ulam sequences and of the finite “interval-with-mask” patterns that arise in rigidity results. In Theorem 3.2 we construct a single $\Delta^{0}_{1}$ predicate $\mathrm{U}(a,b,m)$ such that for all coprime $a<b$ and $m\in\mathbb{N}$,
$\mathrm{U}(a,b,m)\iff m\in U(a,b).$
In Section 2.3 we define a primitive recursive coding of finite patterns and prove (Lemma 2.6) that for each code $c$ with $\mathrm{Pattern}(c)$, the associated set
$\mathrm{PatternSet}_{c}(a,b)=\{m\in\mathbb{N}:\mathrm{InPattern}(c,a,b,m)\}$
is a finite union of linearly parameterised intervals with periodic masks, in the sense used in rigidity results for $U(a,b)$.
Complexity of rigidity.
Using this coding, we analyse the logical form of natural rigidity statements. For a fixed coprime pair $(a,b)$, strong rigidity asserts that $U(a,b)$ eventually coincides with some pattern set $\mathrm{PatternSet}_{c}(a,b)$. In Theorem 4.3 we show that, for each fixed coprime $(a,b)$, the strong rigidity statement
$\exists c\,\exists N\,\forall m\geq N\,\bigl(\mathrm{U}(a,b,m)\leftrightarrow\mathrm{InPattern}(c,a,b,m)\bigr)$
is equivalent to a $\Sigma^{0}_{2}$ sentence of first-order arithmetic. At the family level, the conjectural statement that every coprime pair $(a,b)$ satisfies strong rigidity,
$\forall a\,\forall b\,\bigl(\gcd(a,b)=1\wedge a<b\rightarrow\mathrm{Rig}(a,b)\bigr),$
is shown in Theorem 4.6 to be equivalent to a $\Pi^{0}_{3}$ sentence. (As noted in Remark 1.1, this family-level conjecture is not expected to hold as stated; strong rigidity appears to fail for $U(a,b)$ with $a>1$. However, the complexity classification remains valid regardless of the statement’s truth value.)
Complexity of regularity and simple combinatorial properties.
Let $U(a,b)=\{u_{k}:k\in\mathbb{N}\}$ be enumerated in increasing order and set $g_{k}=u_{k+1}-u_{k}$ for the associated gap sequence. In Theorem 5.3 we show that, for each fixed coprime $(a,b)$, the natural regularity statement
$\exists N\,\exists p\,\forall k\geq N\,(g_{k+p}=g_{k}),$
asserting eventual periodicity of the gap sequence of $U(a,b)$, is equivalent to a $\Sigma^{0}_{2}$ sentence of first-order arithmetic. In Section 5.2 we show that other basic properties, such as “$U(a,b)$ has only finitely many even elements”, can also be expressed by $\Sigma^{0}_{2}$ sentences for each fixed $(a,b)$.
Complexity of density statements.
In Section 5 we study the asymptotic density of $U(a,b)$ via the counting function $C(a,b,n)=|U(a,b)\cap[0,n]|$. For each fixed coprime $(a,b)$ and rational $q\in\mathbb{Q}$, we show in Theorem 5.8 that the upper- and lower-density inequalities
$\overline{d}(U(a,b))\leq q,\qquad\underline{d}(U(a,b))\geq q,$
and the exact-density statement $d(U(a,b))=q$ are all arithmetical and in fact equivalent (for fixed parameters) to $\Pi^{0}_{3}$ sentences of arithmetic. In particular, all of these statements lie low in the arithmetical hierarchy and are absolutely determined by the structure $(\mathbb{N},+,\times)$.
Absoluteness.
Since all of the sentences appearing in the rigidity, regularity and density results above are arithmetical, standard absoluteness results immediately imply that their truth values do not depend on the ambient set-theoretic universe. In Section 6 we show that if $M$ and $N$ are transitive models of $\mathrm{ZFC}$ with the same natural numbers, then for each coprime $(a,b)$ and each rational $q$ the following have the same truth value in $M$ and $N$: strong rigidity and family rigidity; regularity and finiteness-of-evens; and the density statements described above. In particular, these statements cannot be made true in one forcing extension of $\mathrm{ZFC}$ and false in another while preserving $\omega$.
Model-theoretic consequences.
Finally, we initiate a model-theoretic analysis of the expansions
$\mathcal{U}_{a,b}=(\mathbb{N},+,0,1,\mathrm{U}_{a,b})$
of Presburger arithmetic by the predicate naming $U(a,b)$. In Section 7 we show that if strong rigidity holds for $U(a,b)$ (in the sense of Theorem 4.3), then the predicate $\mathrm{U}_{a,b}$ is definable in the reduct $(\mathbb{N},+,0,1)$ and $\mathcal{U}_{a,b}$ is a definitional expansion of Presburger arithmetic (Lemma 7.6). As a consequence (Proposition 7.11), the theory $\mathrm{Th}(\mathcal{U}_{a,b})$ does not interpret $(\mathbb{N},+,\times)$ and remains model-theoretically tame: it is NIP and dp-minimal, and its definable subsets of $\mathbb{N}$ form families of sets of uniformly bounded VC-dimension.
The detailed statements and proofs of these results are given in Sections 37 below. We emphasise that, beyond the $U(1,n)$ family where rigidity results are known, our theorems concern the logical complexity and absoluteness of the corresponding rigidity, regularity, and density schemas for general coprime $(a,b)$; we do not claim these combinatorial statements hold in full generality.
1.6 Organisation of the paper
Section 1.3 together with the technical overview provide informal and formal background for the main results. Section 2 gives precise definitions: Ulam sequences, the arithmetical hierarchy, absoluteness, and the coding of finite combinatorial patterns. Section 3 constructs the uniform $\Delta^{0}_{1}$ predicate $\mathrm{U}(a,b,m)$ and the pattern predicates used in rigidity statements (Theorem 3.2). Section 4 analyses the complexity of strong rigidity, establishing Theorems 4.3 and 4.6. Section 5 treats regularity and density, proving Theorems 5.3 and 5.8. Section 6 draws absoluteness and proof-theoretic consequences. Section 7 develops the model-theoretic analysis of $(\mathbb{N},+,\mathrm{U}_{a,b})$, proving Proposition 7.11 and formulating tameness conjectures. We conclude in Section 8 with open problems.
1.7 Notation and conventions
We collect here the main notational conventions used throughout the paper.
Symbol Meaning
$\mathbb{N},\mathbb{Z},\mathbb{Q}$ The sets of natural, integer, and rational numbers, respectively.
$U(a,b)$ The Ulam sequence with initial terms $a<b$, seen as a subset of $\mathbb{N}$.
$(u_{k})_{k\in\mathbb{N}}$ The increasing enumeration of $U(a,b)$, so that $U(a,b)=\{u_{k}:k\in\mathbb{N}\}$.
$g_{k}$ The gap sequence $g_{k}=u_{k+1}-u_{k}$ associated to $U(a,b)$.
$\mathrm{U}(a,b,m)$ A fixed $\Delta^{0}_{1}$ predicate in arithmetic coding $m\in U(a,b)$ (Theorem 3.2).
$\mathrm{U}_{a,b}$ The unary predicate $m\mapsto\mathrm{U}(a,b,m)$ in the structure $\mathcal{U}_{a,b}$.
$\mathrm{Pattern}(c)$ Primitive recursive predicate asserting that $c\in\mathbb{N}$ encodes a well-formed finite pattern (Section 2.3).
$\mathrm{InPattern}(c,a,b,m)$ Primitive recursive relation expressing that $m$ lies in the pattern determined by $c$ at parameters $(a,b)$.
$\mathrm{PatternSet}_{c}(a,b)$ The set $\{m\in\mathbb{N}:\mathrm{InPattern}(c,a,b,m)\}$ associated to a pattern code $c$.
$B_{\max}(c,a,b)$ A primitive recursive bound with $\mathrm{PatternSet}_{c}(a,b)\subseteq\{m:m\leq B_{\max}(c,a,b)\}$ (Lemma 3.6).
$C(a,b,n)$ The counting function $C(a,b,n)=|U(a,b)\cap[0,n]|$ (Lemma 5.7).
$\overline{d}(X),\underline{d}(X)$ Upper and lower asymptotic densities of $X\subseteq\mathbb{N}$ as in Section 5.
$d(X)$ The (asymptotic) density of $X$ when $\overline{d}(X)=\underline{d}(X)$.
$\mathcal{U}_{a,b}$ The expansion $(\mathbb{N},+,0,1,\mathrm{U}_{a,b})$ of Presburger arithmetic by the predicate naming $U(a,b)$ (Section 7).
$\mathrm{Th}(\mathcal{M})$ The complete first-order theory of a structure $\mathcal{M}$.
Unless otherwise specified, we work in the standard model $(\mathbb{N},+,\times,0,1)$ of first-order arithmetic. We freely identify subsets of $\mathbb{N}$ with their characteristic predicates, and use the usual notation $\Sigma^{0}_{n},\Pi^{0}_{n},\Delta^{0}_{n}$ for the levels of the arithmetical hierarchy.
1.8 Motivations
Beyond the specific complexity bounds established here, one of our main motivations is conceptual. Ulam sequences provide a concrete and well-studied test case in additive combinatorics, and the rigidity phenomena observed by Hinman–Kuca–Schlesinger–Sheydvasser exhibit a striking blend of structure and apparent irregularity. Our results show that a large class of natural questions about this behaviour — including strong rigidity, gap regularity, and density — are in fact of very low logical complexity and are absolutely determined by the arithmetic of $(\mathbb{N},+,\times)$. The expansions $(\mathbb{N},+,\mathrm{U}_{a,b})$ form a new family of test structures for the model theory of expansions of Presburger arithmetic, linking combinatorial rigidity to tameness notions such as NIP and dp-minimality.
2 Preliminaries
2.1 Ulam sequences
We begin by recalling the standard definition of Ulam sequences $U(a,b)$ and their higher-dimensional generalisations.
Definition 2.1.
Let $a,b$ be positive integers with $a<b$. The Ulam sequence $U(a,b)$ is the increasing sequence of positive integers $(u_{k})_{k\in\mathbb{N}}$ defined recursively as follows:
  • $u_{1}=a$, $u_{2}=b$;
  • for $k\geq 3$, we set
    $u_{k}\;=\;\min\bigl\{n>u_{k-1}:\exists!\,1\leq i<j<k\,\bigl(n=u_{i}+u_{j}\bigr)\bigr\}.$
We write $U(a,b)$ for the underlying set $\{u_{k}:k\in\mathbb{N}\}$, and we write $m\in U(a,b)$ if $m$ is one of these terms.
In this paper we will mostly restrict attention to coprime pairs $a<b$; when $\gcd(a,b)>1$ the sequence exhibits obvious degeneracies (for instance, it remains contained in the arithmetic progression $a+\gcd(a,b)\mathbb{N}$). The classical example is $U(1,2)$, introduced by Ulam [Ulam1964], whose initial terms are
$U(1,2)=\{1,2,3,4,6,8,11,13,16,18,26,28,\dots\}.$
More generally, Ulam sequences $U(a,b)$ are special instances of the “$1$-additive” sequences studied by Finch, Schmerl–Spiegel and Cassaigne–Finch [Finch1992ZeroAdditive, SchmerlSpiegel1994, CassaigneFinch1995], where each new term is required to have a unique representation as a sum of two distinct earlier terms.
It is immediate from the definition that each $U(a,b)$ is infinite and strictly increasing, and that the membership relation $m\in U(a,b)$ is decidable by simulating the recursive construction up to $m$. In particular, each $U(a,b)$ is a recursive subset of $\mathbb{N}$. We will exploit this effectivity in Section 3.
Remark 2.2 (Higher-dimensional generalisations).
Ulam-type constructions can be defined in higher dimensions: Kravitz–Steinerberger [KravitzSteinerberger2018] introduced Ulam sets in $\mathbb{Z}^{n}$, and Hinman–Kuca–Schlesinger–Sheydvasser [HinmanKucaSchlesingerSheydvasser2019Involve] established rigidity phenomena for these as well. The arithmetisation methods of this paper extend naturally to such settings, but we focus exclusively on the one-dimensional sequences $U(a,b)$ for concreteness.
2.2 The arithmetical hierarchy and absoluteness
We briefly recall the arithmetical hierarchy and the basic absoluteness facts we will use. Our reference for background material is, for example, Shoenfield [Shoenfield1967] or Kunen [Kunen1980].
We work in the usual first-order language of arithmetic $\mathcal{L}_{\mathrm{PA}}=\{0,1,+,\times,\leq\}$. A formula is bounded (or $\Delta^{0}_{0}$) if every quantifier is of the form $\forall x<t$ or $\exists x<t$, where $t$ is a term not containing $x$. The classes $\Sigma^{0}_{n}$ and $\Pi^{0}_{n}$ of the arithmetical hierarchy are defined inductively as follows:
  • $\Sigma^{0}_{0}=\Pi^{0}_{0}$ is the class of bounded formulas;
  • $\Sigma^{0}_{n+1}$ consists of formulas equivalent (over Peano arithmetic) to formulas of the form $\exists x\,\varphi(x,\bar{y})$ with $\varphi\in\Pi^{0}_{n}$;
  • $\Pi^{0}_{n+1}$ consists of formulas equivalent to formulas of the form $\forall x\,\varphi(x,\bar{y})$ with $\varphi\in\Sigma^{0}_{n}$.
A formula is arithmetical if it is $\Sigma^{0}_{n}$ or $\Pi^{0}_{n}$ for some $n$. A set $X\subseteq\mathbb{N}^{k}$ is said to be $\Sigma^{0}_{n}$ (respectively, $\Pi^{0}_{n}$) if it is definable in $(\mathbb{N},+,\times)$ by a $\Sigma^{0}_{n}$ (respectively, $\Pi^{0}_{n}$) formula; it is $\Delta^{0}_{n}$ if it is both $\Sigma^{0}_{n}$ and $\Pi^{0}_{n}$. In particular, recursive (computable) subsets of $\mathbb{N}^{k}$ are exactly the $\Delta^{0}_{1}$ subsets, and primitive recursive relations are also $\Delta^{0}_{1}$.
Throughout the paper we will freely switch between recursive/primitive recursive definitions and low-level arithmetical formulas: whenever we say that a relation $R(\bar{x})$ is primitive recursive (or recursive), we may equivalently view it as being defined by a fixed $\Delta^{0}_{1}$ formula $R(\bar{x})$ in $\mathcal{L}_{\mathrm{PA}}$.
The complexity classification in the arithmetical hierarchy interacts well with the set-theoretic notion of absoluteness. If $M$ is a transitive model of $\mathrm{ZFC}$, we write $\mathbb{N}^{M}$ for its natural numbers. A sentence $\varphi$ in the language of arithmetic is said to be arithmetical if it is $\Sigma^{0}_{n}$ or $\Pi^{0}_{n}$ for some $n$ (equivalently, if all of its quantifiers range over $\mathbb{N}$ and it uses only $0,1,+,\times,\leq$).
Lemma 2.3 (Absoluteness for arithmetical sentences).
Let $M$ and $N$ be transitive models of $\mathrm{ZFC}$ with the same natural numbers, i.e. $\mathbb{N}^{M}=\mathbb{N}^{N}=\mathbb{N}$. Then for every arithmetical sentence $\varphi$ in the language of arithmetic, we have
$M\models\varphi\quad\Longleftrightarrow\quad N\models\varphi.$
Proof.
The satisfaction relation for arithmetical formulas can be defined by primitive recursion on the complexity of formulas, using only quantification over $\mathbb{N}$. In particular, the truth value of an arithmetical sentence depends only on the underlying structure (,+,×,)($\mathbb{N}$$\times,\leq$). Since $M$ and $N$ have the same natural numbers and the same interpretations of $\times$$\leq,$ they assign the same truth value to every arithmetical sentence. ∎
In particular, if $V$ is the ambient universe and $V[G]$ is a forcing extension, then $V$ and $V[G]$ have the same natural numbers and the same structure (,+,×,)($\mathbb{N}$$\times,\leq$), so they agree on all arithmetical sentences. Thus the truth value of any arithmetical statement is invariant under set forcing. The main complexity results of this paper (for rigidity, regularity and density statements about $U(a,b)$) will show that all of these can be expressed by arithmetical sentences of low complexity, and hence enjoy this strong form of absoluteness.
2.3 Coding finite data by natural numbers
In this subsection we fix, once and for all, a standard coding of finite combinatorial data by natural numbers. This will allow us to treat the “finite patterns” appearing in rigidity statements as single natural numbers, and to work uniformly with a primitive recursive decoding procedure.
We assume familiarity with classical Gödel coding of finite tuples and finite subsets of $\mathbb{N}$. For definiteness, one may proceed as follows.
  • A finite tuple $(x_{0},\dots,x_{k-1})$ of natural numbers is coded by
    $\langle x_{0},\dots,x_{k-1}\rangle\;=\;p_{0}^{x_{0}+1}\cdots p_{k-1}^{x_{k-1}+1},$
    where $(p_{i})_{i\in\mathbb{N}}$ is the increasing sequence of prime numbers. The corresponding decoding functions $\mathrm{len}$ and $\mathrm{entry}(c,i)$, which recover the length and $i$th entry of a coded tuple $c$, are primitive recursive.
  • A finite subset $S\subseteq\{0,\dots,L-1\}$ is coded by the natural number
    $\mathrm{code}_{L}(S)\;=\;\sum_{i\in S}2^{i}.$
    Conversely, given $L$ and a code $s<2^{L}$, membership $i\in S$ is equivalent to the $i$th binary digit of $s$ being $1$, which is again a primitive recursive relation.
We will not need to commit to a specific coding scheme; any fixed primitive recursive encoding of finite tuples and finite subsets will do. What matters is that the encoding and decoding operations are primitive recursive, so that all of the relations we define below are $\Delta^{0}_{1}$.
Rigidity statements for Ulam sequences $U(a,b)$ (and for their higher dimensional analogues) typically express that, beyond some threshold, the set $U(a,b)$ can be described as a finite union of intervals whose endpoints depend linearly on the parameters, together with a periodic “mask” specifying which points in each interval are included. We now formalise this as follows.
Definition 2.4.
A pattern component consists of the following data:
  • integers $A_{1},A_{2},B_{1},B_{2},p,q,L$ with $L\geq 1$;
  • a finite subset $S\subseteq\{0,\dots,L-1\}$.
Given such a component and a pair $(a,b)\in\mathbb{N}^{2}$, we define the associated interval
$I(a,b)=[A(a,b),B(a,b)]\cap\mathbb{N},$
where $A(a,b)=A_{1}a+A_{2}b+p$ and $B(a,b)=B_{1}a+B_{2}b+q$, and the associated masked subset
$C(a,b)=\bigl\{m\in I(a,b):(m-A(a,b))\bmod L\in S\bigr\}.$
A finite pattern for a family $U(a,b)$ is then a finite family of pattern components, which for each $(a,b)$ gives rise to a finite union of masked subsets $C(a,b)$. We encode such patterns by single natural numbers.
Definition 2.5.
A pattern code is a natural number $c$ which, under a fixed primitive recursive decoding scheme, encodes a finite family of pattern components, i.e. a finite list of tuples of the form
$(A_{1},A_{2},B_{1},B_{2},p,q,L,s),$
where $L\geq 1$ and $s$ is the code of a subset $S\subseteq\{0,\dots,L-1\}$. Given such a code $c$ and parameters $(a,b)\in\mathbb{N}^{2}$, the associated pattern set is
$\mathrm{PatternSet}_{c}(a,b)\;=\;\bigcup_{j<k}C_{j}(a,b),$
where $(C_{j}(a,b))_{j<k}$ ranges over the masked subsets determined by the components decoded from $c$.
Formally, we will not work directly with the sets $C_{j}(a,b)$ and $\mathrm{PatternSet}_{c}(a,b)$, but rather with the corresponding membership relations, which we now record.
Lemma 2.6.
There exists a primitive recursive predicate $\mathrm{Pattern}(c)$ expressing that $c$ encodes a well-formed pattern, and a primitive recursive relation $\mathrm{InPattern}(c,a,b,m)$ such that whenever $\mathrm{Pattern}(c)$ holds, we have
$\mathrm{InPattern}(c,a,b,m)\quad\Longleftrightarrow\quad m\in\mathrm{PatternSet}_{c}(a,b).$
In particular, for each fixed $(a,b)$ the set $\mathrm{PatternSet}_{c}(a,b)\subseteq\mathbb{N}$ is $\Delta^{0}_{1}$-definable uniformly in $c$.
Proof.
By construction, a pattern code $c$ consists of a finite list of components, each of which is a finite tuple of integers. Using our fixed coding of finite tuples, we may assume that $c$ is itself a code of such a list. The predicate $\mathrm{Pattern}(c)$ simply asserts, via bounded primitive recursive checks, that $c$ decodes to a list of tuples of the required shape and that each component satisfies the basic well-formedness conditions (for instance, $L\geq 1$ and the mask code $s$ is less than $2^{L}$). All of these checks are primitive recursive in the entries of the decoded tuples.
Given a well-formed code $c$, to decide whether $m\in\mathrm{PatternSet}_{c}(a,b)$ we proceed as follows. We first decode $c$ into its finite list of components. For each component $(A_{1},A_{2},B_{1},B_{2},p,q,L,s)$ we compute $A(a,b)$ and $B(a,b)$ and check whether $m\in[A(a,b),B(a,b)]$. If so, we compute $r=(m-A(a,b))\bmod L$ and test whether the $r$th binary digit of $s$ is $1$, that is, whether $2^{r}$ occurs in the binary expansion of $s$. We then take the disjunction over all components. Since the number of components and all the relevant ranges are bounded by primitive recursive functions of $c,a,b,m$, this procedure defines a primitive recursive relation $\mathrm{InPattern}(c,a,b,m)$.
By construction, whenever $\mathrm{Pattern}(c)$ holds, the above procedure agrees with the intended semantic definition of $\mathrm{PatternSet}_{c}(a,b)$, so the displayed equivalence follows. The final claim about $\Delta^{0}_{1}$ definability is immediate from the observation that primitive recursive relations correspond to $\Delta^{0}_{1}$ formulas in the language of arithmetic. ∎
Remark 2.7 (The family $U(1,n)$: mask-free decompositions).
For the specific family $U(1,n)$, the pattern components considered above specialise to the “piecewise linear” descriptions that have been computed experimentally in the literature. In particular, Sheydvasser has implemented algorithms which, given the numerical data of $U(1,2),\dots,U(1,14)$, recover coefficients $c_{j},d_{j}$ (with $c_{j}\to\infty$ as $j\to\infty$) such that for all sufficiently large $n$ in a fixed residue class modulo some period $L$, one has a decomposition of the form
$U(1,n)\cap[1,c_{j}n+d_{j}]\;=\;\bigcup_{i<k}\bigl([a_{i}n+b_{i},\,a^{\prime}_{i}n+b^{\prime}_{i}]\cap\mathbb{N}\bigr);$
see [Sheydvasser2021LinearPolynomials, SheydvasserUlamCode]. Crucially, for the family $U(1,n)$ there are no masks: every interval $a_{i}n+b_{i},\,a^{\prime}_{i}n+b^{\prime}_{i}]$ contributes all of its integer points to $U(1,n)$, without periodic deletions. In our notation, this corresponds to pattern components with $L=1$ and $S=\{0\}$ (the trivial mask).
This mask-free property is specific to the $U(1,n)$ family. For general $U(a,b)$ with $a>1$, rigidity phenomena are less well understood, and non-trivial masks may in principle appear; see Remark 1.1 below. The general pattern-coding framework of Section 2.3 accommodates both cases.
In the sequel, we will freely identify a pattern code $c$ with the associated family of sets $(\mathrm{PatternSet}_{c}(a,b))_{(a,b)\in\mathbb{N}^{2}}$. In particular, rigidity statements of the form “for all sufficiently large $m$ we have $m\in U(a,b)$ if and only if $m\in\mathrm{PatternSet}_{c}(a,b)$” can be expressed as first-order sentences quantifying only over natural numbers and using the primitive recursive predicates $\mathrm{Pattern}$ and $\mathrm{InPattern}$.
3 Uniform coding of Ulam sequences in arithmetic
3.1 A $\Delta^{0}_{1}$ definition of $\mathrm{U}(a,b,m)$
We first show that membership in $U(a,b)$ is uniformly decidable by a Turing machine which takes $(a,b,m)$ as input. We then invoke the standard fact that recursive relations on $\mathbb{N}$ are $\Delta^{0}_{1}$-definable in $(\mathbb{N},+,\times)$ to obtain a uniform $\Delta^{0}_{1}$ predicate.
Lemma 3.1 (Algorithmic decidability of membership).
There is a Turing machine $M$ which, given as input a triple $(a,b,m)$ of positive integers with $a<b$, halts with output $1$ if $m\in U(a,b)$ and with output 0 otherwise. In particular, the ternary relation
$R(a,b,m)\;:\!\iff\;m\in U(a,b)$
is recursive (computable), uniformly in $(a,b)$.
Proof.
Fix $a<b$ and $m\in\mathbb{N}$. We describe an algorithm which decides whether $m\in U(a,b)$. The algorithm will construct the initial segment of the Ulam sequence $U(a,b)$ up to the point where either $m$ has appeared or it can no longer appear.
Initialisation. Set $S_{2}=\{a,b\}$ and $u_{1}=a$, $u_{2}=b$. If $m\in\{a,b\}$, halt with output $1$.
Inductive step. Suppose we have constructed $S_{k}=\{u_{1},\dots,u_{k}\}$, an initial segment of $U(a,b)$, where $u_{1}<u_{2}<\dots<u_{k}$ and $k\geq 2$. We now search for the next Ulam term $u_{k+1}$ as follows.
For each integer $n$ strictly larger than $u_{k}$, in increasing order, we test whether $n$ has a unique representation as a sum of two distinct elements of $S_{k}$. That is, we count the number of unordered pairs $\{i,j\}\subseteq\{1,\dots,k\}$ with $i\neq j$ such that $u_{i}+u_{j}=n$. This involves only finitely many additions and comparisons, hence is a finite computation. We proceed with $n=u_{k}+1,u_{k}+2,\dots$, stopping at the first $n$ for which there is exactly one such representation.
There are two possibilities:
  • (1)
    If the first such $n$ satisfies $n\leq m$, we set $u_{k+1}=n$ and $S_{k+1}=S_{k}\cup\{n\}$, and repeat the inductive step with $k:=k+1$. If $n=m$, then at this stage we have found $m$ as an element of $U(a,b)$ and we may halt with output $1$.
  • (2)
    If the first such $n$ satisfies $n>m$, then no further Ulam terms are $\leq m$. Indeed, by definition $u_{k+1}$ is the smallest admissible integer larger than $u_{k}$, so any subsequent term $u_{k^{\prime}}$ (for $k^{\prime}>k$) must satisfy $u_{k^{\prime}}\geq u_{k+1}>m$. In particular, no new elements of $U(a,b)\cap[1,m]$ will ever appear beyond this stage. If we reach such an $n>m$ without having seen $m$ among $\{u_{1},\dots,u_{k}\}$, then $m\notin U(a,b)$ and we may halt with output 0.
Since $U(a,b)$ is infinite and strictly increasing, there is always a next Ulam term $u_{k+1}$, and the search over $n>u_{k}$ will eventually find it. Therefore, for each fixed $(a,b,m)$, the above procedure halts after finitely many steps with the correct answer. Moreover, the algorithm is uniform: the code of the Turing machine does not depend on $(a,b)$ or $m$ and operates the same way for all inputs.
This shows that the characteristic function of the relation $R(a,b,m)\iff m\in U(a,b)$ is computable. Hence $R$ is a recursive ternary relation on $\mathbb{N}$, uniformly in the parameters $(a,b)$. ∎
We now translate this into a $\Delta^{0}_{1}$ definition in the language of arithmetic. The argument is standard: recursive relations are exactly the relations definable by $\Delta^{0}_{1}$ formulas in $(\mathbb{N},+,\times)$, via Kleene’s $T$-predicate and the representability of recursive functions in Peano arithmetic (see, for instance, [Shoenfield1967, Chapter VI]).
Theorem 3.2 (Uniform arithmetisation of Ulam sequences).
There exists a $\Delta^{0}_{1}$ formula $\mathrm{U}(a,b,m)$ in the language of arithmetic such that for all coprime $a<b$ and all $m\in\mathbb{N}$, we have
$\mathrm{U}(a,b,m)\iff m\in U(a,b).$
Moreover, the definition is uniform in $(a,b)$: the same formula $\mathrm{U}(a,b,m)$ works for all input parameters $(a,b)$.
Proof.
By Lemma 3.1, the relation
$R(a,b,m)\;:\!\iff\;m\in U(a,b)$
is recursive (computable) as a subset of $\mathbb{N}^{3}$, uniformly in $(a,b)$. It is a standard theorem of recursion theory that every recursive relation on $\mathbb{N}$ is definable in $(\mathbb{N},+,\times)$ by a $\Delta^{0}_{1}$ formula: concretely, one can fix a universal Turing machine, code computation histories by natural numbers, and define a $\Sigma^{0}_{1}$ formula expressing that a computation halts with a given output in a given number of steps; the complement of a recursive relation is also recursively enumerable, so the same relation is $\Pi^{0}_{1}$ as well, hence $\Delta^{0}_{1}$ (see, e.g., [Shoenfield1967, Theorem VI.1.1]).
Applying this to the recursive relation $R(a,b,m)$, we obtain a $\Delta^{0}_{1}$ formula $\mathrm{U}(a,b,m)$ such that
$(\mathbb{N},+,\times)\models\mathrm{U}(a,b,m)\quad\Longleftrightarrow\quad R(a,b,m),$
i.e. $\mathrm{U}(a,b,m)$ holds exactly when $m\in U(a,b)$. The construction of $\mathrm{U}$ depends only on the fixed universal Turing machine and the coding of computation histories, not on the particular values of $(a,b)$, so the definition is uniform in $(a,b)$ as claimed. ∎
3.2 Coding patterns for rigidity
We now package the pattern machinery from Section 2.3 into a form tailored to rigidity statements. Recall that Lemma 2.6 provides a primitive recursive predicate $\mathrm{Pattern}(c)$ and a primitive recursive relation $\mathrm{InPattern}(c,a,b,m)$ indicating membership of $m$ in the pattern determined by $c$ at parameters $(a,b)$.
Definition 3.3.
Given a pattern code $c$ satisfying $\mathrm{Pattern}(c)$, we define the associated pattern set at parameters $(a,b)\in\mathbb{N}^{2}$ by
$\mathrm{PatternSet}_{c}(a,b)\;=\;\{m\in\mathbb{N}:\mathrm{InPattern}(c,a,b,m)\}.$
Thus for each fixed $c$ and $(a,b)$, the set $\mathrm{PatternSet}_{c}(a,b)$ is the finite union of masked intervals $C_{j}(a,b)$ determined by the components encoded in $c$, as described in Section 2.3. We record two elementary facts that will be used repeatedly.
Definition 3.4 (Masked interval).
Let $A,B,L\in\mathbb{Z}$ with $A\leq B$ and $L\geq 1$, and let $S\subseteq\{0,\dots,L-1\}$ be finite. The masked interval determined by $(A,B,L,S)$ is the set
$I(A,B,L,S)=\{m\in\mathbb{N}:A\leq m\leq B\text{ and }(m-A)\bmod L\in S\}.$
Thus $I(A,B,L,\{0,\dots,L-1\})$ is the full interval $A,B]\cap\mathbb{N}$, while proper subsets $S$ encode periodic deletion of residue classes modulo $L$.
Lemma 3.5.
For each pattern code $c$, the relation
$P_{c}(a,b,m)\;:\!\iff\;\mathrm{InPattern}(c,a,b,m)$
is $\Delta^{0}_{1}$ in the structure $(\mathbb{N},+,\times)$; in particular, for each fixed $c$ and $(a,b)$ the set $\mathrm{PatternSet}_{c}(a,b)\subseteq\mathbb{N}$ is recursive.
Proof.
By Lemma 2.6, the relation $\mathrm{InPattern}(c,a,b,m)$ is primitive recursive as a subset of $\mathbb{N}^{4}$, uniformly in $c$. Every primitive recursive relation is recursive, hence definable by a $\Delta^{0}_{1}$ formula in the language of arithmetic (see, for instance, [Shoenfield1967, Chapter VI]). This yields the claim. ∎
Next we isolate an explicit upper bound on the support of a pattern set.
Lemma 3.6.
There exists a primitive recursive function $B_{\max}(c,a,b)$ such that for every pattern code $c$ with $\mathrm{Pattern}(c)$ and every $(a,b)\in\mathbb{N}^{2}$ we have
$\mathrm{PatternSet}_{c}(a,b)\subseteq\{m\in\mathbb{N}:m\leq B_{\max}(c,a,b)\}.$
In particular, for fixed $(c,a,b)$ the set $\mathrm{PatternSet}_{c}(a,b)$ is finite, with size bounded by a primitive recursive function of $(c,a,b)$.
Proof.
Fix a pattern code $c$ with $\mathrm{Pattern}(c)$. By definition, $c$ encodes a finite list of components
$(A_{1}^{(j)},A_{2}^{(j)},B_{1}^{(j)},B_{2}^{(j)},p^{(j)},q^{(j)},L^{(j)},s^{(j)})\quad(j<k),$
each of which contributes a masked subset $C_{j}(a,b)\subseteq\mathbb{N}$ contained in an interval
$I_{j}(a,b)=[A^{(j)}(a,b),B^{(j)}(a,b)]\cap\mathbb{N},$
where $A^{(j)}(a,b)=A_{1}^{(j)}a+A_{2}^{(j)}b+p^{(j)}$ and $B^{(j)}(a,b)=B_{1}^{(j)}a+B_{2}^{(j)}b+q^{(j)}$. The pattern set is
$\mathrm{PatternSet}_{c}(a,b)=\bigcup_{j<k}C_{j}(a,b)\subseteq\bigcup_{j<k}I_{j}(a,b).$
Define
$B_{\max}(c,a,b)=\max_{j<k}B^{(j)}(a,b).$
Since $k$ and the tuples $(A_{1}^{(j)},\dots,q^{(j)})$ are all decoded from $c$ by primitive recursive functions, the function $B_{\max}$ is primitive recursive in $(c,a,b)$. By construction, for every $j<k$ and every $m\in I_{j}(a,b)$, we have $m\leq B^{(j)}(a,b)\leq B_{\max}(c,a,b)$. Hence every $m\in\mathrm{PatternSet}_{c}(a,b)$ satisfies $m\leq B_{\max}(c,a,b)$, as required.
The final statement about finiteness follows immediately, since a subset of $\{0,1,\dots,B_{\max}(c,a,b)\}$ is finite and its size is bounded by $B_{\max}(c,a,b)+1$, which is primitive recursive in $(c,a,b)$. ∎
Later, when we formulate strong rigidity statements, we will typically assert that beyond some threshold $N$ the recursive set $U(a,b)$ coincides with a pattern set $\mathrm{PatternSet}_{c}(a,b)$. Lemma 3.5 ensures that membership in the right-hand side is $\Delta^{0}_{1}$-definable, and Lemma 3.6 allows us to bound the range of $m$ that must be checked when comparing $U(a,b)$ with a given pattern set on an initial segment. Together, these facts will be used in Section 4 to show that the natural rigidity statements occupy low levels of the arithmetical hierarchy.
Example 3.7 (Encoding a single linear block).
As a concrete illustration, consider the family $U(1,n)$ for $n\geq 4$ studied in [HinmanKucaSchlesingerSheydvasser2019JNT, Sheydvasser2021LinearPolynomials]. Table 1 of [Sheydvasser2021LinearPolynomials] (reproduced there as “The first 30 intervals of U(1,X)U(1,X)”) shows, among others, a block of the form
$4X+2,\,5X-1]\subseteq U(1,X),$
which upon evaluation at $X=n$ gives the interval $4n+2,\,5n-1]\subseteq U(1,n)$ for all sufficiently large $n$.
In our notation, this single block can be realised by a single pattern component with coefficients
$A_{1}=0,\quad A_{2}=4,\quad B_{1}=0,\quad B_{2}=5,\quad p=2,\quad q=-1,\quad L=1,\quad S=\{0\}.$
For parameters $(a,b)=(1,n)$ this gives
$A(1,n)=A_{1}\cdot 1+A_{2}\cdot n+p=4n+2,\qquad B(1,n)=B_{1}\cdot 1+B_{2}\cdot n+q=5n-1,$
so the associated contribution to $\mathrm{PatternSet}_{c}(1,n)$ is exactly
$\{m\in\mathbb{N}:4n+2\leq m\leq 5n-1\}.$
since the mask $S=\{0\}$ and period $L=1$ simply mean that we take all points in the interval. More complicated components, with $L>1$ and $S\subsetneq\{0,\dots,L-1\}$, encode periodic deletion of residue classes inside such linearly parameterised intervals; see Section 2.3 for the general scheme.
For instance, taking $L=3$ and $S=\{0,2\}$ would keep only those $m$ in the interval with $(m-(4n+2))\bmod 3\in\{0,2\}$, that is, deleting every third element.
4 Complexity of rigidity statements for Ulam sequences
4.1 Rigidity for a fixed $(a,b)$
We now spell out the strong rigidity statement for a fixed pair $(a,b)$ in the language of arithmetic and analyse its position in the arithmetical hierarchy.
Definition 4.1 (Strong rigidity schema).
Fix coprime $a<b$. The strong rigidity statement $\mathrm{Rig}(a,b)$ asserts that there exist a pattern code $c$ and a cutoff $N$ such that for all $m\geq N$,
$\mathrm{U}(a,b,m)\iff\mathrm{InPattern}(c,a,b,m).$
Equivalently, there exist $c,N\in\mathbb{N}$ such that
$U(a,b)\cap[N,\infty)\;=\;\mathrm{PatternSet}_{c}(a,b)\cap[N,\infty).$
Remark 4.2 (Initial-segment formulation).
For the family $U(1,n)$, rigidity results are often stated in an equivalent “initial-segment” form: there exist coefficients $c_{j},d_{j}$ (with $c_{j}\to\infty$ as $j\to\infty$) such that for all sufficiently large $n$ in a fixed residue class,
$U(1,n)\cap[1,c_{j}n+d_{j}]\;=\;\mathrm{PatternSet}_{c}(1,n)\cap[1,c_{j}n+d_{j}].$
This formulation better reflects how rigidity is computationally verified: one checks that $U(1,n)$ matches the predicted pattern on successively larger initial segments $1,c_{j}n+d_{j}]$. The two formulations are equivalent for our purposes, since the pattern set is eventually equal to the Ulam sequence if and only if they agree on all sufficiently large initial segments.
Recall that by Theorem 3.2, the predicate $\mathrm{U}(a,b,m)$ is $\Delta^{0}_{1}$, uniformly in $(a,b)$, and by Lemma 3.5 the relation $\mathrm{InPattern}(c,a,b,m)$ is also $\Delta^{0}_{1}$, uniformly in $c$ and $(a,b)$. We now show that for each fixed $(a,b)$ the sentence $\mathrm{Rig}(a,b)$ is equivalent (over Peano arithmetic) to a $\Sigma^{0}_{2}$ sentence.
Theorem 4.3.
For each fixed coprime pair $(a,b)$, the statement $\mathrm{Rig}(a,b)$ is equivalent to a $\Sigma^{0}_{2}$ sentence of first-order arithmetic.
Proof.
Fix coprime $a<b$. Working in the language of arithmetic, we view $a$ and $b$ as specific numerals and regard $\mathrm{U}(a,b,m)$ and $\mathrm{InPattern}(c,a,b,m)$ as formulas with free variables $c,N,m$ only.
By definition, $\mathrm{Rig}(a,b)$ asserts the existence of $c,N\in\mathbb{N}$ such that
$\forall m\bigl(m\geq N\rightarrow(\mathrm{U}(a,b,m)\leftrightarrow\mathrm{InPattern}(c,a,b,m))\bigr).$
Thus, as a first-order sentence, $\mathrm{Rig}(a,b)$ has the shape
$\exists c\,\exists N\,\forall m\,\Phi(c,N,m),$
where
$\Phi(c,N,m)\;:=\;\bigl(m<N\bigr)\,\vee\,\bigl(\mathrm{U}(a,b,m)\leftrightarrow\mathrm{InPattern}(c,a,b,m)\bigr).$
We first treat $\mathrm{U}$ and $\mathrm{InPattern}$ as atomic predicates. In that case $\Phi(c,N,m)$ is quantifier-free, and hence the displayed sentence is literally of the form $\exists c\,\exists N\,\forall m\,\psi(c,N,m)$ with $\psi$ quantifier-free. By the usual syntactic definition of the arithmetical hierarchy, this shows that (with $\mathrm{U}$ and $\mathrm{InPattern}$ taken as additional relation symbols) $\mathrm{Rig}(a,b)$ is a $\Sigma^{0}_{2}$ sentence.
To interpret this in pure arithmetic, we recall that both $\mathrm{U}(a,b,m)$ and $\mathrm{InPattern}(c,a,b,m)$ are $\Delta^{0}_{1}$ in $(\mathbb{N},+,\times)$, uniformly in the parameters (Theorem 3.2 and Lemma 3.5). In particular, each of them is equivalent (in Peano arithmetic) to both a $\Sigma^{0}_{1}$ formula and a $\Pi^{0}_{1}$ formula in the language of arithmetic. Let us write $\mathrm{U}_{\Sigma}(a,b,m)$ and $\mathrm{InPattern}_{\Sigma}(c,a,b,m)$ for fixed $\Sigma^{0}_{1}$-formulas defining the same relations, and similarly $\mathrm{U}_{\Pi}$ and $\mathrm{InPattern}_{\Pi}$ for fixed $\Pi^{0}_{1}$ definitions.
Then the equivalence
$\mathrm{U}(a,b,m)\leftrightarrow\mathrm{InPattern}(c,a,b,m)$
is equivalent (over Peano arithmetic) to each of
$\bigl(\mathrm{U}_{\Sigma}(a,b,m)\leftrightarrow\mathrm{InPattern}_{\Pi}(c,a,b,m)\bigr)\quad\text{and}\quad\bigl(\mathrm{U}_{\Pi}(a,b,m)\leftrightarrow\mathrm{InPattern}_{\Sigma}(c,a,b,m)\bigr),$
since all four formulas define the same relation. Each of these new equivalences has the form of a conjunction of a $\Sigma^{0}_{1}$ formula and a $\Pi^{0}_{1}$ formula; in particular, it is equivalent to both a $\Sigma^{0}_{1}$ and a $\Pi^{0}_{1}$ formula (by taking appropriate normal forms). It follows that we may choose a single $\Delta^{0}_{1}$ formula $\Theta(c,N,m)$ in the language of arithmetic which is equivalent to $\Phi(c,N,m)$ for all $c,N,m$.
Substituting $\Theta$ for $\Phi$ in the displayed sentence, we obtain an equivalent sentence of the form
$\exists c\,\exists N\,\forall m\,\Theta(c,N,m),$
with $\Theta$ a $\Delta^{0}_{1}$ formula. By the definition of the arithmetical hierarchy, this is a $\Sigma^{0}_{2}$ sentence. Thus $\mathrm{Rig}(a,b)$ is arithmetically equivalent (over Peano arithmetic) to a $\Sigma^{0}_{2}$ sentence for each fixed coprime pair $(a,b)$, as claimed. ∎
4.2 Family-level rigidity
We now pass from a fixed pair $(a,b)$ to uniform statements about all coprime pairs. For simplicity we quantify over all positive integers $a<b$ and restrict attention to coprime pairs via the condition $\gcd(a,b)=1$.
Definition 4.4 (Family rigidity schema).
The family rigidity statement $\mathrm{FRig}$ asserts that
$\forall a\,\forall b\,\bigl(\gcd(a,b)=1\wedge a<b\rightarrow\mathrm{Rig}(a,b)\bigr),$
where $\mathrm{Rig}(a,b)$ is the strong rigidity statement from the previous subsection.
Remark 4.5.
As discussed in Remark 1.1, the statement $\mathrm{FRig}$ is not expected to be true: computational evidence suggests that $\mathrm{Rig}(a,b)$ fails for many pairs with $a>1$. We analyse the complexity of $\mathrm{FRig}$ nonetheless, both because it illustrates the general method and because understanding why $\mathrm{FRig}$ is $\Pi^{0}_{3}$ clarifies the logical structure of family-level rigidity statements. A more refined family-level statement, restricting to $a=1$, would have the same complexity.
Since the greatest common divisor is a primitive recursive function, the predicate $\gcd(a,b)=1$ is $\Delta^{0}_{1}$ in $(\mathbb{N},+,\times)$. Thus the additional coprimality and ordering conditions do not affect the quantifier alternation pattern in any essential way. We now show that $\mathrm{FRig}$ is a $\Pi^{0}_{3}$ sentence.
Theorem 4.6.
The family rigidity statement $\mathrm{FRig}$ is equivalent to a $\Pi^{0}_{3}$ sentence of first-order arithmetic.
Proof.
By Theorem 4.3, for each fixed coprime pair $(a,b)$ the strong rigidity statement $\mathrm{Rig}(a,b)$ is equivalent (over Peano arithmetic) to a $\Sigma^{0}_{2}$ sentence. Unwinding the definition, we may write this as
$\mathrm{Rig}(a,b)\quad\equiv\quad\exists c\,\exists N\,\forall m\,\Phi(a,b,c,N,m),$
where $\Phi$ is a $\Delta^{0}_{1}$ formula in the language of arithmetic, built from the predicates $\mathrm{U}(a,b,m)$ and $\mathrm{InPattern}(c,a,b,m)$ and bounded comparisons such as $m<N$.
Substituting this into the definition of $\mathrm{FRig}$, we see that $\mathrm{FRig}$ is equivalent to
$\forall a\,\forall b\,\Bigl(\gcd(a,b)=1\wedge a<b\rightarrow\exists c\,\exists N\,\forall m\,\Phi(a,b,c,N,m)\Bigr).$
Using the usual logical equivalence $P\rightarrow Q\equiv\neg P\vee Q$ and distributing the disjunction, we can rewrite the matrix as
$\neg\bigl(\gcd(a,b)=1\wedge a<b\bigr)\ \vee\ \exists c\,\exists N\,\forall m\,\Phi(a,b,c,N,m).$
The condition $\gcd(a,b)=1\wedge a<b$ is $\Delta^{0}_{1}$, and $\Phi$ is $\Delta^{0}_{1}$, so the disjunction of these two clauses is still equivalent to a $\Sigma^{0}_{2}$ formula in the variables $(a,b,c,N,m)$. Concretely, we may treat $\Phi$ as a $\Sigma^{0}_{1}$ formula (since it is $\Delta^{0}_{1}$) and observe that $\neg(\gcd(a,b)=1\wedge a<b)$ is $\Pi^{0}_{0}$; their disjunction is then arithmetically equivalent to a $\Sigma^{0}_{1}$ formula, and the outer $\exists c\,\exists N\,\forall m$ gives a $\Sigma^{0}_{2}$ matrix in $(a,b)$.
Thus $\mathrm{FRig}$ is equivalent to a sentence of the form
$\forall a\,\forall b\,\Psi(a,b),$
where $\Psi(a,b)$ is a $\Sigma^{0}_{2}$ formula of arithmetic. Equivalently, by coding pairs $(a,b)$ as single natural numbers and reindexing, we may express $\mathrm{FRig}$ in prenex form as
$\forall n\,\exists c\,\exists N\,\forall m\,\Theta(n,c,N,m),$
where $\Theta$ is $\Delta^{0}_{1}$. This has quantifier prefix $\forall\exists\exists\forall$, with three alternating blocks (universal, existential, universal), and hence is a $\Pi^{0}_{3}$ sentence in the arithmetical hierarchy.
Therefore the family rigidity statement $\mathrm{FRig}$ is equivalent to a $\Pi^{0}_{3}$ sentence of first-order arithmetic. ∎
Similar arguments apply to more refined versions of family-level rigidity, for instance those in which the pattern code $c$ is allowed to depend only on the congruence class of $b$ modulo some fixed integer $L$, or in which the endpoints of the intervals in the pattern are required to be linear forms in $(a,b)$ with prescribed coefficients. In each case the additional requirements are expressed by $\Delta^{0}_{1}$ conditions on the parameters and codes, and so do not increase the arithmetical complexity beyond $\Pi^{0}_{3}$.
5 Complexity of regularity and density statements
5.1 Eventual periodicity of gaps
We begin by formalising the notion of regularity via the gap sequence.
Definition 5.1.
Let $(u_{k})_{k\in\mathbb{N}}$ be the increasing enumeration of $U(a,b)$. The gap sequence associated to $U(a,b)$ is the sequence
$g_{k}=u_{k+1}-u_{k}\qquad(k\in\mathbb{N}).$
We say that $U(a,b)$ is regular if there exist $N,p\in\mathbb{N}$ such that for all $k\geq N$ we have $g_{k+p}=g_{k}$, i.e.
$\exists N\,\exists p\,\forall k\geq N\,\bigl(u_{k+1}-u_{k}=u_{k+p+1}-u_{k+p}\bigr).$
To analyse the logical complexity of this statement, we need to know that the enumeration $k\mapsto u_{k}$ is itself recursive (and hence $\Delta^{0}_{1}$-definable) uniformly in $(a,b)$. This is immediate from the algorithmic construction of $U(a,b)$.
Lemma 5.2 (Recursive enumeration of $U(a,b)$).
There is a Turing machine $M^{\prime}$ which, given as input a triple $(a,b,k)$ of positive integers with $a<b$, halts with output $u_{k}$, where $(u_{k})$ is the increasing enumeration of $U(a,b)$. In particular, the function
$E(a,b,k)=u_{k}$
is a total recursive function of $(a,b,k)$. Consequently, there is a $\Delta^{0}_{1}$ formula $\mathrm{UTerm}(a,b,k,m)$ in the language of arithmetic such that for all $(a,b,k,m)\in\mathbb{N}^{4}$,
$\mathrm{UTerm}(a,b,k,m)\quad\Longleftrightarrow\quad m=u_{k}.$
Proof.
Fix $a<b$ and $k\in\mathbb{N}$. We describe an algorithm which computes $u_{k}$. The construction is essentially the same as in the proof of Lemma 3.1 but now we stop when we reach the $k$th term rather than a fixed bound $m$.
Initialisation. Set $S_{2}=\{a,b\}$ and $u_{1}=a$, $u_{2}=b$. If $k=1$ (respectively, $k=2$), halt with output $a$ (respectively, $b$).
Inductive step. Suppose $k\geq 3$ and we have constructed $S_{j}=\{u_{1},\dots,u_{j}\}$ for some $j\geq 2$, where $u_{1}<\dots<u_{j}$ are the first $j$ terms of $U(a,b)$. To construct $u_{j+1}$, we search over integers $n>u_{j}$ in increasing order and count the number of unordered pairs $\{i,\ell\}\subseteq\{1,\dots,j\}$ with $i\neq\ell$ and $u_{i}+u_{\ell}=n$. We stop at the first $n$ for which there is exactly one such representation, and declare $u_{j+1}=n$. We then set $S_{j+1}=S_{j}\cup\{n\}$ and continue.
This procedure constructs the Ulam sequence $U(a,b)$ term by term. Since $U(a,b)$ is infinite and strictly increasing, the process continues indefinitely, and for each input $k$ the algorithm reaches the stage $j=k$ after finitely many steps, at which point $u_{k}$ has been computed. Thus the function $E(a,b,k)=u_{k}$ is computable (recursive) and uniform in the parameters $(a,b,k)$.
As in the proof of Theorem 3.2, every total recursive function on $\mathbb{N}^{3}$ is representable in $(\mathbb{N},+,\times)$ by a $\Delta^{0}_{1}$ formula: there is a $\Delta^{0}_{1}$ formula $\mathrm{UTerm}(a,b,k,m)$ such that $\mathrm{UTerm}(a,b,k,m)$ holds if and only if $m$ is the output of the fixed Turing machine $M^{\prime}$ on input $(a,b,k)$; see, for instance, [Shoenfield1967, Theorem VI.1.1]. By construction, this output is $u_{k}$, so $\mathrm{UTerm}(a,b,k,m)\leftrightarrow m=u_{k}$ as required. ∎
We can now express regularity entirely in terms of the $\Delta^{0}_{1}$ predicate $\mathrm{UTerm}$.
Theorem 5.3.
For each fixed coprime $(a,b)$, the statement
$\exists N\,\exists p\,\forall k\geq N\,\bigl(g_{k+p}=g_{k}\bigr)$
expressing regularity of $U(a,b)$ is equivalent to a $\Sigma^{0}_{2}$ sentence of first-order arithmetic.
Proof.
Fix coprime $a<b$. Working in the language of arithmetic, we treat $a$ and $b$ as parameters and use the $\Delta^{0}_{1}$ formula $\mathrm{UTerm}(a,b,k,m)$ from Lemma 5.2 to refer to the $k$th Ulam term.
The condition $g_{k+p}=g_{k}$ can be rewritten as
$u_{k+1}-u_{k}\;=\;u_{k+p+1}-u_{k+p}.$
In terms of $\mathrm{UTerm}$ this says that for all quadruples $(m_{0},m_{1},n_{0},n_{1})\in\mathbb{N}^{4}$ with
$\mathrm{UTerm}(a,b,k,m_{0}),\quad\mathrm{UTerm}(a,b,k+1,m_{1}),\quad\mathrm{UTerm}(a,b,k+p,n_{0}),\quad\mathrm{UTerm}(a,b,k+p+1,n_{1}),$
we have $m_{1}-m_{0}=n_{1}-n_{0}$. Since the $u_{\ell}$ are uniquely determined by $(a,b,\ell)$, there is in fact exactly one quadruple of natural numbers $(m_{0},m_{1},n_{0},n_{1})$ satisfying these four instances of $\mathrm{UTerm}$, but we do not need to use uniqueness explicitly.
Thus the regularity of $U(a,b)$ is equivalent (over Peano arithmetic) to the sentence
$\exists N\,\exists p\,\forall k\geq N\,\forall m_{0}\,\forall m_{1}\,\forall n_{0}\,\forall n_{1}\,\Bigl(\bigwedge_{\begin{subarray}{c}\ell\in\{k,k+1,k+p,k+p+1\}\\ x\in\{m_{0},m_{1},n_{0},n_{1}\}\end{subarray}}\mathrm{UTerm}(\dots)\ \rightarrow\ m_{1}-m_{0}=n_{1}-n_{0}\Bigr),$
where the conjunction $\bigwedge\mathrm{UTerm}(\dots)$ abbreviates the four instances
$\mathrm{UTerm}(a,b,k,m_{0})\ \wedge\ \mathrm{UTerm}(a,b,k+1,m_{1})\ \wedge\ \mathrm{UTerm}(a,b,k+p,n_{0})\ \wedge\ \mathrm{UTerm}(a,b,k+p+1,n_{1}).$
More explicitly, we can write this in prenex form as
$\exists N\,\exists p\,\forall k\,\forall m_{0}\,\forall m_{1}\,\forall n_{0}\,\forall n_{1}\,\Psi(a,b,N,p,k,m_{0},m_{1},n_{0},n_{1}),$
where $\Psi$ is the matrix
$\Psi:=\bigl(k<N\bigr)\ \vee\ \Bigl(\bigwedge_{\begin{subarray}{c}\ell\in\{k,k+1,k+p,k+p+1\}\\ x\in\{m_{0},m_{1},n_{0},n_{1}\}\end{subarray}}\mathrm{UTerm}(\dots)\ \rightarrow\ m_{1}-m_{0}=n_{1}-n_{0}\Bigr).$
The condition $k<N$ is bounded and hence $\Delta^{0}_{0}$. Each instance of $\mathrm{UTerm}$ is $\Delta^{0}_{1}$ by Lemma 5.2, and the arithmetic equation $m_{1}-m_{0}=n_{1}-n_{0}$ is quantifier-free. Therefore $\Psi$ is a $\Delta^{0}_{1}$ formula in the free variables $(a,b,N,p,k,m_{0},m_{1},n_{0},n_{1})$. The overall sentence has quantifier prefix
$\exists N\,\exists p\,\forall k\,\forall m_{0}\,\forall m_{1}\,\forall n_{0}\,\forall n_{1},$
which consists of one existential block followed by one universal block. By the definition of the arithmetical hierarchy, this is a $\Sigma^{0}_{2}$ sentence.
Thus the regularity of $U(a,b)$ is arithmetically equivalent to a $\Sigma^{0}_{2}$ sentence for each fixed coprime pair $(a,b)$, as claimed. ∎
5.2 Finiteness of even terms and related properties
We next consider simple combinatorial properties of $U(a,b)$ such as having only finitely many even elements. These also sit low in the arithmetical hierarchy once expressed in terms of the predicate $\mathrm{U}(a,b,m)$.
Proposition 5.4.
For fixed coprime $(a,b)$, the statement $U(a,b)$ has only finitely many even elements” is equivalent to a $\Sigma^{0}_{2}$ sentence of arithmetic.
Proof.
Fix coprime $a<b$. The set of even elements of $U(a,b)$ is
$E(a,b)=\{m\in\mathbb{N}:\text{$m$ is even and }\mathrm{U}(a,b,m)\}.$
Saying that $E(a,b)$ is finite is equivalent to saying that there exists an $N$ such that no even $m\geq N$ belongs to $U(a,b)$, i.e.
$\exists N\,\forall m\,\Bigl(m\geq N\wedge 2\mid m\ \rightarrow\ \neg\mathrm{U}(a,b,m)\Bigr).$
Call this sentence $(\ast)$.
First, if $(\ast)$ holds, then there is some $N$ such that every even element of $U(a,b)$ is less than $N$, so $E(a,b)\subseteq\{0,2,4,\dots,N-2\}$ and is in particular finite. Conversely, if $E(a,b)$ is finite, there is some $N$ greater than all elements of $E(a,b)$. Then by definition no even $m\geq N$ lies in $U(a,b)$, so $(\ast)$ holds. Thus $(\ast)$ is equivalent (over Peano arithmetic) to the informal statement that $U(a,b)$ has only finitely many even elements.
We now analyse the logical form of $(\ast)$. The outer quantifiers are $\exists N\,\forall m$. The matrix
$m\geq N\wedge 2\mid m\ \rightarrow\ \neg\mathrm{U}(a,b,m)$
is built from bounded inequalities, the divisibility relation $2\mid m$ and the predicate $\mathrm{U}(a,b,m)$. The inequality $m\geq N$ and the divisibility condition $2\mid m$ are $\Delta^{0}_{0}$ formulas, and $\mathrm{U}(a,b,m)$ is $\Delta^{0}_{1}$ by Theorem 3.2. Hence the whole matrix is $\Delta^{0}_{1}$. Therefore $(\ast)$ has the form
$\exists N\,\forall m\,\Phi(N,m),$
with $\Phi$ a $\Delta^{0}_{1}$ formula. By the definition of the arithmetical hierarchy, this is a $\Sigma^{0}_{2}$ sentence. This completes the proof. ∎
Remark 5.5.
The negation of the finiteness statement, namely $U(a,b)$ has infinitely many even elements”, can be written as
$\forall N\,\exists m\geq N\,\bigl(2\mid m\wedge\mathrm{U}(a,b,m)\bigr),$
which has the quantifier prefix $\forall\exists$ with a $\Delta^{0}_{1}$ matrix and hence is $\Pi^{0}_{2}$. Similar reasoning applies to other simple properties such as $U(a,b)$ contains only finitely many terms in a given residue class modulo $\ell$ or $U(a,b)$ contains infinitely many terms in a given residue class modulo $\ell$, which are again $\Sigma^{0}_{2}$ and $\Pi^{0}_{2}$ respectively for each fixed $(a,b)$ and $\ell$.
5.3 Density
We recall the usual notions of upper and lower asymptotic density.
Definition 5.6.
For a set $X\subseteq\mathbb{N}$, the upper and lower asymptotic densities are
$\overline{d}(X)=\limsup_{n\to\infty}\frac{|X\cap[0,n]|}{n+1},\qquad\underline{d}(X)=\liminf_{n\to\infty}\frac{|X\cap[0,n]|}{n+1}.$
If these coincide, we write $d(X)$ for the common value.
For $U(a,b)$ we will work with the associated counting function
$C_{a,b}(n)\;=\;|U(a,b)\cap[0,n]|.$
Since membership in $U(a,b)$ is recursive, so is $C_{a,b}(n)$.
Lemma 5.7 (Counting function).
There is a total recursive function $C(a,b,n)$ such that for all $(a,b,n)\in\mathbb{N}^{3}$,
$C(a,b,n)=|U(a,b)\cap[0,n]|.$
Consequently, there is a $\Delta^{0}_{1}$ formula $\mathrm{Count}(a,b,n,t)$ in the language of arithmetic such that
$\mathrm{Count}(a,b,n,t)\quad\Longleftrightarrow\quad t=C(a,b,n).$
Proof.
Fix $a<b$ and $n\in\mathbb{N}$. To compute $|U(a,b)\cap[0,n]|$, we may proceed as follows. Using the algorithm from Lemma 3.1, we can effectively decide for each $m\leq n$ whether $m\in U(a,b)$, by constructing $U(a,b)$ up to the stage at which either we see $m$ or the next Ulam term exceeds $m$. Running this decision procedure for all $m=0,1,\dots,n$ and incrementing a counter whenever $m\in U(a,b)$ yields the desired cardinality.
This gives a Turing machine which, on input $(a,b,n)$, halts with output $|U(a,b)\cap[0,n]|$. Hence the function $C(a,b,n)$ is recursive. As in the proof of Theorem 3.2, every total recursive function on $\mathbb{N}^{3}$ is representable in $(\mathbb{N},+,\times)$ by a $\Delta^{0}_{1}$ formula: there is a $\Delta^{0}_{1}$ formula $\mathrm{Count}(a,b,n,t)$ expressing that $t$ is the output of the fixed Turing machine computing $C$ on input $(a,b,n)$ (see [Shoenfield1967, Theorem VI.1.1]). By construction, this output is $|U(a,b)\cap[0,n]|$, which yields the equivalence.
We note that the naive algorithm described above computes $C(a,b,n)$ by constructing $U(a,b)$ up to $n$ once and counting membership along the way; in particular, $C$ is not only total recursive but primitive recursive, and its computation can be bounded by an elementary function of $n$ for each fixed $(a,b)$. We do not attempt to optimise the running time here, as only recursiveness is needed for our logical applications. ∎
We now analyse the logical form of the natural density inequalities for $U(a,b)$. Throughout, we regard the rational parameter $q\in\mathbb{Q}$ as fixed, say $q=p/s$ with $p\in\mathbb{Z}$ and $s\in\mathbb{N}$, and we suppress the dependence of primitive recursive bounds on this representation.
Theorem 5.8.
For each fixed coprime $(a,b)$ and each rational $q\in\mathbb{Q}$, the following statements are arithmetical:
  • (i)
    $\overline{d}(U(a,b))\leq q$;
  • (ii)
    $\underline{d}(U(a,b))\geq q$;
  • (iii)
    $d(U(a,b))=q$.
More precisely, for each fixed $(a,b)$ and $q$, the statements (i) and (ii) can be written as $\Pi^{0}_{3}$ sentences of arithmetic, and (iii) is equivalent to the conjunction of (i) and (ii), and hence is again $\Pi^{0}_{3}$. In particular, all three statements are arithmetical and lie (non-optimally) in $\Delta^{0}_{4}$.
Proof.
Fix coprime $(a,b)$ and rational $q\in\mathbb{Q}$. Let $C(n)=C_{a,b}(n)$ for brevity. By Lemma 5.7, the relation $C(n)=t$ is $\Delta^{0}_{1}$, uniformly in $n,t$ and the parameters $(a,b)$.
(i) Upper density $\overline{d}(U(a,b))\leq q$. By definition, $\overline{d}(U(a,b))\leq q$ means that
$\limsup_{n\to\infty}\frac{C(n)}{n+1}\leq q.$
Equivalently, for every $k\in\mathbb{N}_{>0}$ there exists $N\in\mathbb{N}$ such that for all $n\geq N$ we have
$\frac{C(n)}{n+1}\leq q+\frac{1}{k}.$
For a fixed rational $q$, the inequality
$\frac{C(n)}{n+1}\leq q+\frac{1}{k}$
can be rewritten as an inequality between integers using only additions and multiplications. Concretely, if $q=p/s$ with $s\geq 1$, then
$\frac{C(n)}{n+1}\leq\frac{p}{s}+\frac{1}{k}\quad\Longleftrightarrow\quad sk\,C(n)\;\leq\;\bigl(pk+s(n+1)\bigr)\,(n+1),$
and the right-hand side involves only the parameters $(a,b)$ (through $C(n)$), the fixed rational $q$, and the quantified variables $k,n$. Since $C(n)$ is a total recursive function, both sides of this inequality are values of primitive recursive functions of $(a,b,q,k,n)$, and so the relation
$\frac{C(n)}{n+1}\leq q+\frac{1}{k}$
is $\Delta^{0}_{1}$ as a predicate in $k,n$ (with $(a,b,q)$ as parameters).
Thus the statement $\overline{d}(U(a,b))\leq q$” is equivalent to the sentence
$\forall k\,\exists N\,\forall n\geq N\,\Phi(k,N,n),$
where $\Phi(k,N,n)$ is the $\Delta^{0}_{1}$ formula expressing the above inequality. By the definition of the arithmetical hierarchy, this is a $\Pi^{0}_{3}$ sentence.
(ii) Lower density $\underline{d}(U(a,b))\geq q$. By definition, $\underline{d}(U(a,b))\geq q$ means that
$\liminf_{n\to\infty}\frac{C(n)}{n+1}\geq q.$
Equivalently, for every $k\in\mathbb{N}_{>0}$ there exists $N\in\mathbb{N}$ such that for all $n\geq N$ we have
$\frac{C(n)}{n+1}\geq q-\frac{1}{k}.$
As before, for fixed rational $q$ this inequality can be rewritten as an inequality between values of primitive recursive functions of $k,n$ (with parameters $(a,b,q)$), and hence is expressible by a $\Delta^{0}_{1}$ formula $\Psi(k,N,n)$ in the language of arithmetic. This gives the sentence
$\forall k\,\exists N\,\forall n\geq N\,\Psi(k,N,n),$
which again has quantifier prefix $\forall\exists\forall$ with a $\Delta^{0}_{1}$ matrix. Thus it is a $\Pi^{0}_{3}$ sentence.
(iii) Exact density $d(U(a,b))=q$. By definition, $d(U(a,b))=q$ holds if and only if
$\overline{d}(U(a,b))\leq q\quad\text{and}\quad\underline{d}(U(a,b))\geq q.$
By parts (i) and (ii), each of these is equivalent to a $\Pi^{0}_{3}$ sentence in the language of arithmetic (with parameters $(a,b,q)$). Their conjunction is therefore again equivalent to a $\Pi^{0}_{3}$ sentence.
In particular, all three statements (i)–(iii) are arithmetical. Since any $\Pi^{0}_{3}$ sentence can trivially be rewritten as both a $\Sigma^{0}_{4}$ and a $\Pi^{0}_{4}$ sentence by adding dummy quantifiers, each of these statements also lies (non-optimally) in $\Delta^{0}_{4}$ in the sense that it is equivalent to both a $\Sigma^{0}_{4}$ and a $\Pi^{0}_{4}$ sentence. ∎
Combinatorial statement For fixed $(a,b)$ Uniform / family version
Strong rigidity $\mathrm{Rig}(a,b)$ $\Sigma^{0}_{2}$ (Thm. 4.3) $\Pi^{0}_{3}$ (Thm. 4.6)
Gap regularity (eventual periodicity) $\Sigma^{0}_{2}$ (Thm. 5.3)
Finitely many even elements $\Sigma^{0}_{2}$ (Prop. 5.4)
Upper/lower density
$\overline{d}(U(a,b))\leq q$, $\underline{d}(U(a,b))\geq q$ $\Pi^{0}_{3}$ (Thm. 5.8)
Exact density $d(U(a,b))=q$ $\Pi^{0}_{3}$ (Thm. 5.8)
Figure 1: Logical complexity for rigidity, regularity and density statements.
6 Absoluteness and proof-theoretic strength
6.1 Absoluteness for arithmetical sentences
We briefly recall the general principle underlying the absoluteness results we obtain for rigidity, regularity and density statements. Let $M$ and $N$ be transitive models of $\mathrm{ZFC}$, and write $\mathbb{N}^{M}$ and $\mathbb{N}^{N}$ for their respective sets of natural numbers. If $\varphi$ is a sentence in the language of arithmetic $\mathcal{L}_{\mathrm{PA}}=\{0,1,+,\times,\leq\}$ which is arithmetical (i.e. built using only first-order quantification over $\mathbb{N}$ and the arithmetic symbols), then its truth value depends only on the structure (,+,×,)($\mathbb{N}$$\times,\leq$).
In particular, if $M$ and $N$ have the same natural numbers, i.e. $\mathbb{N}^{M}=\mathbb{N}^{N}=\mathbb{N}$, then they agree on the truth of all arithmetical sentences.
This is exactly Lemma 2.3 from Section 2, and its proof rests on the fact that the satisfaction relation for arithmetical formulas is primitive recursive in the underlying structure (,+,×,)($\mathbb{N}$$\times,\leq$); see, for instance, [Shoenfield1967, Chapter VI]; for a related discussion of definability and absoluteness in set theory, see [Kunen1980, Chapter IV and its appendices].
In particular, if $V$ is the ambient universe and $V[G]$ is any forcing extension, then $V$ and $V[G]$ have the same natural numbers and the same structure (,+,×,)($\mathbb{N}$$\times,\leq$), so they agree on all arithmetical sentences.
6.2 Absoluteness of rigidity and regularity statements
We now apply Lemma 2.3 to the specific statements analysed in Sections 4 and 5. The key input is that all of these statements are arithmetical and in fact occupy low levels of the arithmetical hierarchy.
Corollary 6.1.
The following statements are arithmetical and hence absolute between transitive models of $\mathrm{ZFC}$ with the same natural numbers:
  • for each fixed coprime $(a,b)$, the strong rigidity statement $\mathrm{Rig}(a,b)$;
  • the family rigidity statement $\mathrm{FRig}$;
  • for each fixed coprime $(a,b)$, the regularity statement asserting eventual periodicity of the gap sequence of $U(a,b)$;
  • for each fixed coprime $(a,b)$ and each rational $q\in\mathbb{Q}$, the density statements from Theorem 5.8.
Proof.
For fixed $(a,b)$, Theorem 4.3 shows that $\mathrm{Rig}(a,b)$ is equivalent to a $\Sigma^{0}_{2}$ sentence of arithmetic. Theorem 4.6 shows that $\mathrm{FRig}$ is equivalent to a $\Pi^{0}_{3}$ sentence. Theorem 5.3 shows that regularity of $U(a,b)$ is equivalent to a $\Sigma^{0}_{2}$ sentence. Finally, Theorem 5.8 shows that the upper and lower density inequalities, and hence the exact density statements, are equivalent to $\Pi^{0}_{3}$ sentences for fixed $(a,b)$ and rational $q$.
In particular, in each case the statement in question is arithmetical (i.e. $\Sigma^{0}_{n}$ or $\Pi^{0}_{n}$ for some $n$). Lemma 2.3 therefore implies that any two transitive models of $\mathrm{ZFC}$ with the same natural numbers agree on the truth value of each of these statements. ∎
Remark 6.2.
The transitivity assumption in the preceding corollary can be relaxed. If $M$ and $N$ are $\omega$-models of $\mathrm{ZFC}$ (i.e. their natural numbers are isomorphic to the standard $\mathbb{N}$), then they agree on all arithmetical sentences, regardless of transitivity. In particular, the rigidity, regularity and density statements considered in this paper have the same truth value in any two $\omega$-models of $\mathrm{ZFC}$, and hence are insensitive to variations in the higher set-theoretic universe even in that more general sense.
Corollary 6.3.
None of the above statements can be made true in one forcing extension of $\mathrm{ZFC}$ and false in another forcing extension that preserves $\omega$. In particular, their truth values cannot depend on the Continuum Hypothesis, forcing axioms, or large cardinal axioms.
Proof.
Let $\varphi$ be any of the statements listed in the preceding corollary. Suppose $V$ is a model of $\mathrm{ZFC}$ and $V[G]$ is a forcing extension obtained by set forcing. Then $V$ and $V[G]$ are both transitive models of $\mathrm{ZFC}$ with the same natural numbers: forcing does not create new finite ordinals, so $\omega^{V}=\omega^{V[G]}$ and (,+,×,)($\mathbb{N}$$\times,\leq$) is the same structure in both models. By the previous corollary, $V$ and $V[G]$ must agree on the truth value of $\varphi$. Thus $\varphi$ is neither forced nor refuted by any such forcing.
The same argument applies if we enlarge $\mathrm{ZFC}$ by additional axioms such as $\mathrm{CH}$, forcing axioms or large cardinal hypotheses: any transitive model $M$ of an extended theory $\mathrm{ZFC}+\Gamma$ has a well-defined natural numbers object $\mathbb{N}^{M}$, and any two such models with the same $\mathbb{N}$ agree on all the arithmetical sentences considered above. In particular, no such axiom can change the truth value of $\varphi$. ∎
6.3 Reverse-mathematical remarks
We conclude this section with a brief comment on the reverse-mathematical strength of our complexity analyses. A detailed study would take us too far afield, but the basic point is that the classification of our Ulam statements as low-level arithmetical sentences already holds over weak base theories.
Throughout this subsection we work in the framework of second-order arithmetic as presented in Simpson [Simpson2009]. Recall that $\mathsf{RCA}_{0}$ is the standard base system consisting of recursive comprehension and $\Sigma^{0}_{1}$-induction.
Proposition 6.4.
Over $\mathsf{RCA}_{0}$, the following results are provable:
  • for each fixed coprime $(a,b)$, the strong rigidity statement $\mathrm{Rig}(a,b)$ is equivalent to a $\Sigma^{0}_{2}$ sentence of first-order arithmetic;
  • the family rigidity statement $\mathrm{FRig}$ is equivalent to a $\Pi^{0}_{3}$ sentence;
  • for each fixed coprime $(a,b)$, the regularity statement for $U(a,b)$ is equivalent to a $\Sigma^{0}_{2}$ sentence;
  • for each fixed coprime $(a,b)$ and rational $q\in\mathbb{Q}$, the density statements in Theorem 5.8 are equivalent to $\Pi^{0}_{3}$ sentences.
Sketch of proof.
The proofs of the complexity results in Sections 45 rely only on: (i) Gödel coding of finite tuples and finite sets; (ii) the formalisation of primitive recursive functions and relations; and (iii) the representability of recursive relations by $\Delta^{0}_{1}$ formulas in $(\mathbb{N},+,\times)$, together with routine manipulations of quantifier alternations.
All of these ingredients are available in $\mathsf{RCA}_{0}$. In particular, Chapter II of [Simpson2009] develops Gödel coding of finite sets and sequences in $\mathsf{RCA}_{0}$, while Section III.1 shows that primitive recursive functions and relations are formalised there and that every primitive recursive relation is $\Delta^{0}_{1}$-definable over $(\mathbb{N},+,\times)$. The definition of the arithmetical hierarchy and the usual normal-form theorems for $\Sigma^{0}_{n}$- and $\Pi^{0}_{n}$-formulas can also be carried out in $\mathsf{RCA}_{0}$.
Thus the arguments of Theorems 4.3, 4.6, 5.3, and 5.8 can be formalised in $\mathsf{RCA}_{0}$ without change, yielding the stated equivalences. ∎
In particular, if any of the strong rigidity, regularity, or density statements is true in the standard model $\mathbb{N}$, then its truth is already captured by relatively weak subsystems of second-order arithmetic: no set- existence axioms beyond those of $\mathsf{RCA}_{0}$ are required to prove that the statement is equivalent to a low-level arithmetical sentence. A natural direction for further work would be to investigate the exact reverse- mathematical strength of global assertions such as “every coprime $(a,b)$ satisfies $\mathrm{Rig}(a,b)$” or suitable uniform regularity statements, in the spirit of the calibrations in [Simpson2009].
7 Model-theoretic aspects of $(\mathbb{N},+,\mathrm{U}_{a,b})$
7.1 Model-theoretic background
We briefly recall the model-theoretic notions that appear in Section 7. We work in the usual framework of first-order model theory; see, for example, standard texts on stability and NIP theories for further background.
Definition 7.1 (Independence property (IP) and NIP).
Let $T$ be a complete first-order theory and $\varphi(x;y)$ a formula in the language of $T$, where $x$ and $y$ are tuples of variables. We say that $\varphi(x;y)$ has the independence property (IP) if there exists a model $M\models T$ and a sequence of parameters $(b_{i})_{i\in\mathbb{N}}$ from $M^{|y|}$ such that for every subset $S\subseteq\mathbb{N}$ there is $a_{S}\in M^{|x|}$ with
$M\models\varphi(a_{S};b_{i})\quad\Longleftrightarrow\quad i\in S\qquad\text{for all }i\in\mathbb{N}.$
In words, the family of definable sets $\{\varphi(M;b_{i}):i\in\mathbb{N}\}$ realises all patterns on $\mathbb{N}$.
A theory $T$ is said to be NIP (“not the independence property”) if no formula $\varphi(x;y)$ has IP in any model of $T$. Equivalently, every $\varphi(x;y)$ has finite Vapnik–Chervonenkis dimension in the sense recalled below.
Definition 7.2 (VC-dimension).
Let $X$ be a set and $\mathcal{F}\subseteq\mathcal{P}(X)$ a family of subsets of $X$. A finite subset $A\subseteq X$ is said to be shattered by $\mathcal{F}$ if
$\{F\cap A:F\in\mathcal{F}\}=\mathcal{P}(A),$
i.e. every subset of $A$ arises as the intersection of $A$ with some $F\in\mathcal{F}$. The VC-dimension of $\mathcal{F}$, written $\mathrm{VCdim}(\mathcal{F})$, is the supremum of the cardinalities $|A|$ of finite sets $A\subseteq X$ that are shattered by $\mathcal{F}$ (or $\infty$ if no such finite bound exists).
Given a formula $\varphi(x;y)$ and a structure $M$, we consider the family
$\mathcal{F}_{\varphi}=\{\varphi(M;b):b\in M^{|y|}\}$
of subsets of $M^{|x|}$. We say that $\varphi$ has finite VC-dimension in $M$ if $\mathrm{VCdim}(\mathcal{F}_{\varphi})<\infty$, and we say that $T$ is NIP iff every formula has finite VC-dimension in every model of $T$.
Definition 7.3 (dp-minimality).
A complete NIP theory $T$ is called dp-minimal if, informally, every definable one-parameter family behaves like a “linear” family from the point of view of dependence. There are several equivalent formal definitions; we sketch one that suffices for our purposes.
We say that $T$ is dp-minimal if for every model $M\models T$ and every formula $\varphi(x;y)$ with $|x|=1$, one cannot find two sequences of parameters $(b_{i}^{0})_{i\in\mathbb{N}}$ and $(b_{j}^{1})_{j\in\mathbb{N}}$ from $M^{|y|}$ and elements $(a_{i,j})_{i,j\in\mathbb{N}}$ of $M$ such that
$M\models\varphi(a_{i,j};b_{k}^{\ell})\quad\Longleftrightarrow\quad\bigl(\ell=0\ \wedge\ k=i\bigr)\text{ or }\bigl(\ell=1\ \wedge\ k=j\bigr)$
for all $i,j,k\in\mathbb{N}$ and $\ell\in\{0,1\}$. In other words, there is no “independent pattern” of depth $2$ in a single variable. For more systematic treatments and equivalent characterisations (via dp-rank, VC-density, etc.), one may consult the standard literature on NIP and dp-minimal theories.
A basic example is Presburger arithmetic: the complete theory $\mathrm{Th}(\mathbb{N},+,0,1)$ of addition on the natural numbers is NIP and dp-minimal, and its definable sets are semilinear.
Definition 7.4 (Interpretation and definitional expansion).
Let $\mathcal{M}$ and $\mathcal{N}$ be first-order structures. We say that $\mathcal{N}$ is interpretable in $\mathcal{M}$ if there is a definable set $D\subseteq M^{k}$ (for some $k$) together with definable relations and functions on $D$ making $(D,\dots)$ isomorphic to $\mathcal{N}$. We say that a theory $T$ interprets a structure $\mathcal{N}$ if some (or equivalently every) model of $T$ does.
An expansion $\mathcal{M}^{\prime}$ of a structure $\mathcal{M}$ is called a definitional expansion if every new symbol in the language of $\mathcal{M}^{\prime}$ is definable in $\mathcal{M}$, and conversely every $\mathcal{M}^{\prime}$-formula is equivalent (in $\mathrm{Th}(\mathcal{M})$) to a formula in the original language. Definitional expansions preserve model-theoretic properties such as NIP, dp-minimality, and interpretability power.
For general background on NIP theories, VC-dimension, and dp-minimality we refer to Simon’s monograph [SimonNIP].
7.2 The structure $(\mathbb{N},+,\mathrm{U}_{a,b})$
For each fixed coprime pair $(a,b)$ we consider the expansion of Presburger arithmetic by a unary predicate naming the Ulam sequence $U(a,b)$.
Definition 7.5.
For fixed coprime $(a,b)$ we write
$\mathcal{U}_{a,b}=(\mathbb{N},+,0,1,\mathrm{U}_{a,b}),$
where $\mathrm{U}_{a,b}(m)$ abbreviates $\mathrm{U}(a,b,m)$, the $\Delta^{0}_{1}$ predicate from Theorem 3.2 defining membership in $U(a,b)$.
The reduct of $\mathcal{U}_{a,b}$ to the language $\{0,1,+\}$ is the standard model of Presburger arithmetic, which admits quantifier elimination and whose definable subsets of $\mathbb{N}^{k}$ are exactly the semilinear sets, i.e. finite unions of linear sets and singletons. In particular, every definable subset of $\mathbb{N}$ is a finite union of arithmetic progressions and isolated points. The expansion by $\mathrm{U}_{a,b}$ need not preserve this semilinear behaviour in general: for an arbitrary recursive subset $P\subseteq\mathbb{N}$, the structure $(\mathbb{N},+,0,1,P)$ may define sets far more complicated than semilinear ones, and may even interpret multiplication.
However, if $U(a,b)$ exhibits strong combinatorial rigidity in the sense of Section 4, then the new predicate $\mathrm{U}_{a,b}$ turns out to be definable in the underlying Presburger structure. In that case $\mathcal{U}_{a,b}$ is in fact a definitional expansion of Presburger arithmetic, and inherits its model-theoretic tameness.
We first record the definability of $U(a,b)$ under a strong rigidity assumption.
Lemma 7.6 (Definability of $U(a,b)$ under strong rigidity).
Fix coprime $(a,b)$ and suppose that $\mathrm{Rig}(a,b)$ holds, witnessed by a pattern code $c$ and a cutoff $N$, in the sense that
$\forall m\geq N\,\bigl(\mathrm{U}(a,b,m)\leftrightarrow\mathrm{InPattern}(c,a,b,m)\bigr).$
Then $U(a,b)$ is definable in the Presburger structure $(\mathbb{N},+,0,1)$ with parameters $a,b,c,N$. In particular, there is a formula $\varphi(x)$ in the language $\{0,1,+\}$ with parameters from $\mathbb{N}$ such that for all $m\in\mathbb{N}$,
$\mathcal{U}_{a,b}\models\mathrm{U}_{a,b}(m)\quad\Longleftrightarrow\quad(\mathbb{N},+,0,1)\models\varphi(m).$
Proof.
By assumption there exist $c,N\in\mathbb{N}$ such that for all $m\geq N$ we have
$m\in U(a,b)\quad\Longleftrightarrow\quad m\in\mathrm{PatternSet}_{c}(a,b).$
Thus
$U(a,b)\;=\;\bigl(U(a,b)\cap[0,N-1]\bigr)\ \cup\ \bigl(\mathrm{PatternSet}_{c}(a,b)\cap[N,\infty)\bigr).$
The finite set $U(a,b)\cap[0,N-1]$ is definable in $(\mathbb{N},+,0,1)$ with parameters, since it is a finite subset of $\mathbb{N}$ and we may simply list its elements as constants in a disjunction.
It therefore suffices to show that, for fixed $(a,b,c)$, the tail $\mathrm{PatternSet}_{c}(a,b)\cap[N,\infty)$ is Presburger-definable. By definition of pattern codes (Section 2.3), the set $\mathrm{PatternSet}_{c}(a,b)$ is a finite union of masked intervals
$\bigl\{m\in\mathbb{N}:A(a,b)\leq m\leq B(a,b)\text{ and }(m-A(a,b))\bmod L\in S\bigr\},$
where $A(a,b)$ and $B(a,b)$ are fixed integers depending linearly on $a,b$, $L\geq 1$ is a fixed integer, and $S$ is a fixed finite subset of $\{0,\dots,L-1\}$, all recovered from the code $c$. For fixed $(a,b,c)$ these quantities are just natural number parameters.
Each such masked interval can be written as a finite union of segments of arithmetic progressions. Indeed, for each $r\in S$ the condition $(m-A(a,b))\bmod L=r$ with $A(a,b)\leq m\leq B(a,b)$ is equivalent to
$m=A(a,b)+r+L\cdot t\quad\text{for some }t\in\mathbb{N}\text{ with }A(a,b)\leq m\leq B(a,b).$
For fixed constants $A,B,L,r$, the set
$\{m\in\mathbb{N}:m=A+r+Lt\text{ for some }t\}$
is an arithmetic progression, and its intersection with $[A,B]$ is a finite segment of that progression. It is well-known (and straightforward to verify) that such sets are definable in $(\mathbb{N},+,0,1)$: multiplication by a fixed constant $L$ is definable by repeated addition, and the existence of $t$ with $m=A+r+Lt$ can be expressed by a first-order formula in the language $\{0,1,+\}$ with the constant $L$ as a parameter.
More explicitly, for fixed integers $A,B,L,r$ the set
$X_{A,B,L,r}=\{m\in\mathbb{N}:A\leq m\leq B\text{ and }(m-A)\bmod L=r\}$
can be defined in $(\mathbb{N},+,0,1)$ by the formula
$\exists t\,\bigl(m=A+r+L\cdot t\ \wedge\ A\leq m\leq B\bigr),$
where $L\cdot t$ abbreviates the $L$-fold sum of $t$ with itself. Thus each masked interval appearing in $\mathrm{PatternSet}_{c}(a,b)$ is a finite union of such $X_{A,B,L,r}$, and hence Presburger-definable with parameters $(a,b,c)$.
Since finite unions and intersections of Presburger-definable sets are again Presburger-definable, it follows that each masked interval, and hence each $\mathrm{PatternSet}_{c}(a,b)$, is definable in $(\mathbb{N},+,0,1)$ with parameters $(a,b,c)$. Intersecting with the definable tail condition $m\geq N$ preserves definability. Therefore the entire set $U(a,b)$ is definable in $(\mathbb{N},+,0,1)$ with parameters $(a,b,c,N)$, as claimed. ∎
In particular, under the hypothesis of Lemma 7.6, the predicate $\mathrm{U}_{a,b}(x)$ is definable in the reduct $(\mathbb{N},+,0,1)$, so $\mathcal{U}_{a,b}$ is a definitional expansion of Presburger arithmetic.
7.3 Rigidity and tameness
We now develop the model-theoretic consequences of various structural conditions on Ulam sequences. We begin by introducing a hierarchy of combinatorial assumptions, ranging from strong rigidity to weaker density conditions.
Definition 7.7 (Hierarchy of structural conditions).
Let $(a,b)$ be coprime with $a<b$. We consider the following conditions on $U(a,b)$, listed roughly from strongest to weakest:
  1. (R1)
    Strong rigidity: There exist $c,N$ such that for all $m\geq N$, $m\in U(a,b)$ iff $\mathrm{InPattern}(c,a,b,m)$.
  2. (R2)
    Gap regularity: The gap sequence $(g_{k})$ is eventually periodic, i.e. $\exists N\,\exists p\,\forall k\geq N\,(g_{k+p}=g_{k})$.
  3. (R3)
    Bounded gaps: There exists $B$ such that $g_{k}\leq B$ for all $k$.
  4. (R4)
    Density existence: The limit $d(U(a,b))=\lim_{n\to\infty}|U(a,b)\cap[0,n]|/n$ exists.
  5. (R5)
    Positive lower density: $\underline{d}(U(a,b))=\liminf_{n\to\infty}|U(a,b)\cap[0,n]|/n>0$.
The implications among these conditions are as follows:
$\text{(R1)}\Rightarrow\text{(R2)}\Rightarrow\text{(R3)}\Rightarrow\text{(R5)},\qquad\text{(R2)}\Rightarrow\text{(R4)}.$
Verification of implications.
[…]
$|U(a,b)\cap[0,n]|\geq\frac{n-a}{B}+1,$
giving $\underline{d}(U(a,b))\geq 1/B>0$. ∎
Remark 7.8.
Note that (R3) does not imply (R4) in general: bounded gaps ensure positive lower density but do not by themselves guarantee that the limit $\lim_{n\to\infty}|U(a,b)\cap[0,n]|/n$ exists. Similarly, (R4) does not imply (R5): there exist infinite sets with asymptotic density zero (e.g. the squares or the primes). It is not known whether any Ulam sequence has density zero; all computed examples suggest positive density, but a proof that $d(U(a,b))>0$ whenever it exists remains open.
We now establish the key consequences of gap regularity.
Proposition 7.9 (Consequences of gap regularity).
Suppose $U(a,b)$ satisfies condition (R2): the gap sequence $(g_{k})$ is eventually periodic with period $p$ and threshold $N$. Then:
  1. (i)
    The density $d(U(a,b))$ exists and equals
    $d(U(a,b))=\frac{p}{\sum_{j=0}^{p-1}g_{N+j}}.$
  2. (ii)
    The set $U(a,b)$ is the union of a finite set and a set that is Presburger-definable with parameters.
  3. (iii)
    The expansion $\mathcal{U}_{a,b}$ is NIP.
Proof.
(i) For $k\geq N$, write $k=N+mp+r$ with $0\leq r<p$ and $m\geq 0$. The partial sums of gaps satisfy
$u_{N+mp+r}=u_{N}+m\cdot G+\sum_{j=0}^{r-1}g_{N+j},$
where $G=\sum_{j=0}^{p-1}g_{N+j}$ is the sum of one period of gaps. For large $n$, the number of Ulam elements in $u_{N},n]$ is asymptotically $p\cdot(n-u_{N})/G$. It follows that
$\lim_{n\to\infty}\frac{|U(a,b)\cap[0,n]|}{n}=\frac{p}{G}.$
(ii) The tail $\{u_{k}:k\geq N\}$ consists of elements of the form
$u_{N}+mG+\sum_{j=0}^{r-1}g_{N+j}$
for $m\geq 0$ and $0\leq r<p$. For each fixed $r$, this is an arithmetic progression with first term $u_{N}+\sum_{j=0}^{r-1}g_{N+j}$ and common difference $G$. Thus the tail is a union of $p$ arithmetic progressions. Arithmetic progressions are Presburger-definable, so the tail is Presburger-definable with parameters $(u_{N},g_{N},\dots,g_{N+p-1})$. The finite initial segment $\{u_{k}:k<N\}$ is also Presburger-definable (as a finite set of constants).
(iii) By (ii), there is a formula $\psi(x,\bar{c})$ in the language $\{0,1,+\}$ with a finite tuple of parameters $\bar{c}$ such that $m\in U(a,b)$ iff $(\mathbb{N},+,0,1)\models\psi(m,\bar{c})$. Naming the parameters $\bar{c}$ gives a (definitional) expansion of Presburger arithmetic that remains NIP (since NIP is preserved under naming constants). Now adding a new unary predicate symbol $P(x)$ together with the defining axiom $\forall x\,(P(x)\leftrightarrow\psi(x,\bar{c}))$ is again a definitional expansion, so the resulting theory is NIP. Interpreting $P$ as $\mathrm{U}_{a,b}$ yields $\mathcal{U}_{a,b}$, which is therefore NIP. ∎
Remark 7.10 (Gap regularity vs. strong rigidity).
Lemma 7.6 shows that strong rigidity (R1) implies Presburger-definability of $U(a,b)$ (with parameters), and hence the tameness of $\mathcal{U}_{a,b}$ discussed below. Part (ii) of Proposition 7.9 shows that one actually needs only gap regularity (R2) to obtain Presburger-definability of the tail and therefore of $U(a,b)$ up to a finite set. In particular, the hypothesis of (R2) in Proposition 7.11 below is strictly weaker than the strong rigidity assumption used in Lemma 7.6.
The extra content of strong rigidity is effectivity: it provides an explicit pattern code $c$ from which the Presburger definition of $U(a,b)$ can be computed, whereas gap regularity merely asserts the existence of an eventual periodicity pattern without specifying it in terms of interval-with-mask decompositions. In principle, (R2) is strictly weaker than (R1): one could imagine a sequence with eventually periodic gaps but where the positions of the Ulam elements within each period vary in a way not captured by a simple pattern code. In practice, all known gap-regular examples also satisfy strong rigidity, and it remains open whether (R2) can hold without (R1).
We now record the main model-theoretic consequence. Recall that definitional expansions do not change the class of definable sets (up to naming additional definable predicates) or the interpretability power of a theory: any structure interpretable in a definitional expansion is already interpretable in the original structure.
Proposition 7.11 (Model-theoretic tameness under gap regularity).
Assume that $U(a,b)$ satisfies gap regularity (R2) of Definition 7.7, i.e., its gap sequence $(g_{k})$ is eventually periodic. Then the theory $\mathrm{Th}(\mathcal{U}_{a,b})$ is a definitional expansion of Presburger arithmetic. In particular:
  • $\mathrm{Th}(\mathcal{U}_{a,b})$ does not interpret the full arithmetic structure $(\mathbb{N},+,\times)$;
  • the expansion $\mathcal{U}_{a,b}$ is NIP and dp-minimal; in particular, its definable families of subsets of $\mathbb{N}$ have uniformly bounded VC-dimension.
Proof.
By Proposition 7.9(ii), under the assumption of gap regularity (R2) there is a formula $\varphi(x)$ in the language $\{0,1,+\}$ with parameters such that for all $m\in\mathbb{N}$,
$\mathcal{U}_{a,b}\models\mathrm{U}_{a,b}(m)\quad\Longleftrightarrow\quad(\mathbb{N},+,0,1)\models\varphi(m).$
In other words, $U(a,b)$ is definable (up to a finite initial segment, which is also definable) in the underlying Presburger structure. Thus the predicate symbol $\mathrm{U}_{a,b}$ is definable in the reduct $(\mathbb{N},+,0,1)$, and $\mathcal{U}_{a,b}$ is a definitional expansion of $(\mathbb{N},+,0,1)$: every formula in the expanded language can be translated into an equivalent formula in the original Presburger language by replacing atomic instances of $\mathrm{U}_{a,b}(x)$ with $\varphi(x)$.
Let $T_{\mathrm{Pres}}$ denote the complete theory of $(\mathbb{N},+,0,1)$ and let $T_{U}$ denote $\mathrm{Th}(\mathcal{U}_{a,b})$. Since $T_{U}$ is a conservative definitional extension of $T_{\mathrm{Pres}}$, any structure interpretable in $T_{U}$ is already interpretable in $T_{\mathrm{Pres}}$.
It is a classical fact that Presburger arithmetic does not interpret the full ring $(\mathbb{N},+,\times)$; in particular, multiplication is not interpretable in $(\mathbb{N},+,0,1)$. If $(\mathbb{N},+,\times)$ were interpretable in $T_{U}$, then by definitionality it would be interpretable in $T_{\mathrm{Pres}}$, a contradiction. This proves that $T_{U}$ does not interpret $(\mathbb{N},+,\times)$.
Finally, Presburger arithmetic is NIP and dp-minimal, and NIP and dp-minimality are preserved under definitional expansions. Hence $\mathcal{U}_{a,b}$ is again NIP and dp-minimal, and, in particular, every family of definable subsets of $\mathbb{N}$ in $\mathcal{U}_{a,b}$ has uniformly bounded VC-dimension. Since Presburger arithmetic is decidable, the same holds for $\mathrm{Th}(\mathcal{U}_{a,b})$. ∎
In Section 7.4 we ask how far one can push such tameness without assuming strong rigidity or even gap regularity.
Remark 7.12 (Computational consequences under rigidity).
Under the hypotheses of Proposition 7.11, every first-order sentence in the language of $\mathcal{U}_{a,b}$ can be effectively translated into an equivalent sentence of Presburger arithmetic, uniformly in the parameters $(a,b)$ and in the pattern code witnessing rigidity. In particular, the theory $\mathrm{Th}(\mathcal{U}_{a,b})$ is decidable in this regime, and one may in principle combine standard decision procedures for Presburger arithmetic (such as Cooper’s algorithm or automata-theoretic methods) with the arithmetisation of $U(a,b)$ to obtain explicit algorithms for answering first-order questions about $U(a,b)$. We do not attempt to optimise such algorithms here, as our primary focus is on logical complexity rather than computational complexity.
There is a substantial body of work on decidable and undecidable expansions of Presburger arithmetic; see, for example, Point’s study of decidable extensions [Point2000DecidablePresburger] and Bès’ survey on arithmetical definability [Bes2001SurveyArithDef].
7.4 Unconditional and partial results
The preceding results establish model-theoretic tameness of $\mathcal{U}_{a,b}$ under the hypothesis of gap regularity (R2), and in particular under strong rigidity (R1). In this subsection we investigate what can be said without these assumptions, either unconditionally or under weaker structural hypotheses. While full tameness results remain out of reach in the general case, we identify several constraints that the Ulam construction imposes on the expansion $\mathcal{U}_{a,b}$ and formulate precise conjectures about intermediate cases.
7.4.1 The unique-sum property and its consequences
The defining feature of Ulam sequences—that each term beyond the initial pair has a unique representation as a sum of two distinct earlier terms—is a strong structural constraint. We begin by formalising this and extracting some unconditional consequences.
Definition 7.13 (Representation function).
For a set $P\subseteq\mathbb{N}$ and $n\in\mathbb{N}$, the representation function $r_{P}(n)$ counts the number of ways to write $n$ as a sum of two distinct elements of $P$:
$r_{P}(n)=\bigl|\{(x,y)\in P^{2}:x<y\text{ and }x+y=n\}\bigr|.$
Lemma 7.14 (Representation bound for Ulam sequences).
Let $(a,b)$ be coprime with $a<b$. For every $m\in U(a,b)\setminus\{a,b\}$ we have $r_{U(a,b)}(m)=1$. Moreover, for every $n\notin U(a,b)$ with $n>b$, either $r_{U(a,b)\cap[1,n)}(n)=0$ or $r_{U(a,b)\cap[1,n)}(n)\geq 2$.
Proof.
The first claim is immediate from the definition of the Ulam sequence: an integer $m>b$ is adjoined to $U(a,b)$ at stage $k$ if and only if $m$ is the smallest integer exceeding all previously included terms and having exactly one representation as a sum of two distinct earlier terms. Thus $r_{U(a,b)\cap[1,m)}(m)=1$ at the moment of inclusion, and since no further terms less than $m$ are added subsequently, this remains $r_{U(a,b)}(m)=1$ in the final sequence.
For the second claim, suppose $n>b$ and $n\notin U(a,b)$. At the stage when the Ulam construction considers $n$, either $n$ has no representations as a sum of two distinct earlier terms (so $r_{U(a,b)\cap[1,n)}(n)=0$), or $n$ has at least two such representations (so $r_{U(a,b)\cap[1,n)}(n)\geq 2$). In the latter case, since the Ulam sequence $U(a,b)\cap[1,n)$ is already determined by stage $n$, this representation count persists. ∎
The representation bound suggests heuristic constraints on the density of Ulam sequences, though rigorous upper bounds appear difficult to establish.
Remark 7.15 (Density heuristics).
Intuitively, one expects the density of any Ulam sequence to be bounded well away from $1$. The reason is as follows: if $U=U(a,b)$ had density $d$ close to $1$, then the number of pairs from $U\cap[1,n]$ would be $\binom{dn}{2}\approx d^{2}n^{2}/2$, while the range of possible sums is only $[a+b,2n]$, an interval of length <2n<2n. By the pigeonhole principle, many integers would have multiple representations, and by Lemma 7.14, such integers cannot belong to $U$ (beyond the initial terms). This suggests a tension between high density and the unique-representation requirement.
Making this precise is nontrivial. A naive application of the Cauchy–Schwarz inequality to the second moment $\sum_{m}r_{U_{n}}(m)^{2}$ yields only weak bounds on the number of multiply-represented integers, insufficient to establish a sharp density constraint such as $\overline{d}(U)\leq 1/2$. The difficulty lies in controlling the distribution of representation counts among non-members: while each must have 0 or $\geq 2$ representations, the total “energy” absorbed by non-members depends on how representation counts are distributed, and crude bounds lose too much information.
Numerical evidence suggests that densities of Ulam sequences, when they exist, are quite small. For the classical sequence $U(1,2)$, computations indicate $d(U(1,2))\approx 0.07$ [Steinerberger2017HiddenSignal], and other well-studied pairs yield similarly small values. The gap between these empirical densities and even a conjectural bound of $1/2$ suggests that the rigidity phenomena impose much stronger constraints than naive counting arguments capture.
7.4.2 The additive character of the Ulam construction
We now turn to structural features of the expansion $\mathcal{U}_{a,b}$ that follow from the purely additive nature of the Ulam construction.
Definition 7.16 (Additively closed construction).
A subset $P\subseteq\mathbb{N}$ is said to arise from an additively closed construction if there is a recursive procedure defining $P$ in which membership of $n$ in $P$ depends only on:
  • the values of $P\cap[1,n)$ (the initial segment);
  • linear combinations and comparisons of elements of $P\cap[1,n)$;
  • bounded quantification over $P\cap[1,n)$.
In particular, the procedure never references multiplication, division, exponentiation, or other non-additive operations on the elements themselves.
Lemma 7.17.
For every coprime $(a,b)$, the Ulam sequence $U(a,b)$ arises from an additively closed construction in the sense of Definition 7.16.
Proof.
The Ulam construction proceeds as follows: given $U(a,b)\cap[1,n)$, an integer $n$ is adjoined if and only if (i) $n>\max(U(a,b)\cap[1,n))$, and (ii) there is exactly one pair $(x,y)$ with $\in U(a,b)\cap[1,n), x<yx<y$, and $x+y=n$. Condition (i) involves only comparison. Condition (ii) involves only addition of elements and counting (bounded quantification). No multiplicative operations appear. Thus the construction is additively closed. ∎
The significance of this observation is that additively closed constructions produce predicates whose interaction with the additive structure of $\mathbb{N}$ is constrained. While we cannot prove unconditionally that $\mathcal{U}_{a,b}$ fails to interpret multiplication, we can articulate why such an interpretation would be surprising.
Proposition 7.18 (Decidability of the atomic diagram).
Let $(a,b)$ be coprime with $a<b$. The atomic diagram of $\mathcal{U}_{a,b}=(\mathbb{N},+,0,1,\mathrm{U}_{a,b})$ is decidable: given any quantifier-free sentence $\varphi$ with parameters from $\mathbb{N}$, one can effectively determine whether $\mathcal{U}_{a,b}\models\varphi$.
Proof.
The predicate $\mathrm{U}(a,b,m)$ is $\Delta^{0}_{1}$ by Theorem 3.2, hence decidable. Any quantifier-free formula in the language $\{0,1,+,\mathrm{U}_{a,b}\}$ with parameters involves only finitely many membership queries to $U(a,b)$ and finitely many linear comparisons, all of which are decidable. ∎
Remark 7.19 (On Presburger definability).
Presburger-definable subsets of $\mathbb{N}$ are exactly the eventually periodic sets: finite unions of arithmetic progressions and singletons. It is expected, in line with computational evidence and the conjectured lack of eventual periodicity of gaps for most pairs $(a,b)$, that $\mathrm{U}_{a,b}$ is not definable in $(\mathbb{N},+,0,1)$ without parameters unless strong rigidity holds. Indeed, definability of $U(a,b)$ in Presburger arithmetic would force eventual periodicity of its characteristic function, which appears to fail for generic pairs based on extensive numerical data.
Remark 7.20 (On quantifier elimination).
It seems unlikely that $\mathcal{U}_{a,b}$ admits quantifier elimination in the language $\{0,1,+,\mathrm{U}_{a,b}\}$ when $U(a,b)$ is not eventually periodic. Already the set $\{(x,y)\in\mathbb{N}^{2}:x+y\in U(a,b)\}$ is definable by a simple existential formula, but describing it quantifier-free would require the predicate $\mathrm{U}_{a,b}$ to interact with addition in a highly structured way. We do not pursue a definitive result on quantifier elimination here.
7.4.3 Comparison with known interpretations of arithmetic
To place the model-theoretic status of $\mathcal{U}_{a,b}$ in context, we briefly recall when an expansion $(\mathbb{N},+,P)$ interprets full arithmetic.
Definition 7.21 (Multiplicative encoding).
We say that a predicate $P\subseteq\mathbb{N}$ encodes multiplication if there exist first-order formulas $\varphi_{\times}(x,y,z)$ and $\varphi_{D}(x)$ in the language $\{0,1,+,P\}$ such that $\varphi_{D}$ defines an infinite subset $D\subseteq\mathbb{N}$ and
$(\mathbb{N},+,P)\models\varphi_{\times}(a,b,c)\quad\Longleftrightarrow\quad a\cdot b=c$
for all $a,b,c\in D$. Equivalently, $(\mathbb{N},+,P)$ interprets a nontrivial fragment of $(\mathbb{N},+,\times)$.
Predicates defined in terms of strong multiplicative or exponential growth are known to produce undecidable expansions where arithmetic is interpretable. Classical examples include:
  • $P=\{k^{n}:n\in\mathbb{N}\}$ for a fixed base $k\geq 2$ (powers);
  • $P=\{n!:n\in\mathbb{N}\}$ (factorials);
  • $P=\{p_{n}:n\in\mathbb{N}\}$ (the sequence of primes).
For these and related results, see Villemaire [Villemaire1992], Muchnik–Semenov [MuchnikSemenov2003], Point [Point2000DecidablePresburger], and the survey by Bès [Bes2001SurveyArithDef]. The common feature of these examples is that the predicate $P$ encodes multiplicative information through its growth pattern or divisibility structure.
In contrast, “additively natural” predicates tend to yield tamer expansions.
Remark 7.22 (Sparse sets and tameness; cf. [MuchnikSemenov2003, Bes2001SurveyArithDef]).
Let $P\subseteq\mathbb{N}$ be a set of density zero satisfying suitable sparseness conditions (for example, $|P\cap[0,n]|=O(\log n)$). Under certain additional hypotheses on the growth rate and structure of $P$, the expansion $(\mathbb{N},+,P)$ remains decidable and does not interpret $(\mathbb{N},+,\times)$. The precise conditions vary across different results in the literature; we refer to [MuchnikSemenov2003, Bes2001SurveyArithDef] for details.
Ulam sequences occupy an intermediate regime: they have positive density (conjecturally), so the extreme sparseness results do not directly apply, but they lack the multiplicative structure of the examples above. This motivates the following conjecture.
Conjecture 7.23 (Non-interpretability of multiplication).
For every coprime $(a,b)$, the expansion $\mathcal{U}_{a,b}$ does not interpret $(\mathbb{N},+,\times)$. That is, there is no formula $\varphi(x,y,z)$ in the language $\{0,1,+,\mathrm{U}_{a,b}\}$ defining the graph of multiplication on an infinite domain.
Evidence for this conjecture includes:
  • The purely additive character of the Ulam construction (Lemma 7.17);
  • The absence of any apparent mechanism by which the unique-sum property could encode multiplicative relations;
  • The contrast with known multiplication-encoding predicates, all of which involve exponential, factorial, or similar growth patterns foreign to Ulam sequences.
7.4.4 Weaker structural conditions
We now consider what can be said under conditions weaker than gap regularity (R2). Recall from Definition 7.7 that bounded gaps (R3), density existence (R4), and positive lower density (R5) form progressively weaker assumptions.
For the bounded-gaps case (R3), we expect tameness properties to hold but do not have complete proofs.
Conjecture 7.24 (Tameness under bounded gaps).
Suppose $U(a,b)$ satisfies condition (R3): there exists $B$ such that all gaps $g_{k}$ are at most $B$. Then:
  1. (i)
    The expansion $\mathcal{U}_{a,b}$ is NIP.
  2. (ii)
    For any formula $\varphi(x;y)$ in the language $\{0,1,+,\mathrm{U}_{a,b}\}$ with $|x|=1$, the VC-dimension of the family $\{\varphi(\mathbb{N};b):b\in\mathbb{N}^{|y|}\}$ is finite.
Remark 7.25.
Intuitively, bounded gaps should force some form of tameness: the predicate $\mathrm{U}_{a,b}$ stays within bounded distance of a union of arithmetic progressions (in the sense that between any two consecutive Ulam elements there are at most B1B-1 non-Ulam elements), and one expects the resulting definable families to have bounded VC-dimension. Results of Aschenbrenner–Dolich–Haskell–Macpherson–Starchenko [Aschenbrenner2016VapnikChervonenkis] on VC-density in NIP theories suggest that such bounds should exist, but verifying the hypotheses of their theorems for $\mathcal{U}_{a,b}$ under (R3) alone requires additional work that we do not undertake here.
We summarise the known and conjectured implications in the following table.
Condition NIP dp-minimal No $(\mathbb{N},+,\times)$
Strong rigidity (R1) Yes (Prop. 7.11) Yes (Prop. 7.11) Yes (Prop. 7.11)
Gap regularity (R2) Yes (Props. 7.97.11) Yes (Prop. 7.11) Yes (Prop. 7.11)
Bounded gaps (R3) Conjectured Unknown Conjectured
Density exists (R4) Unknown Unknown Conjectured
Positive lower density (R5) Unknown Unknown Conjectured
No assumption Conjectured Unknown Conjectured
7.4.5 A dichotomy conjecture
The preceding analysis suggests a possible dichotomy for expansions of Presburger arithmetic by Ulam predicates.
Conjecture 7.26 (Ulam dichotomy).
For every coprime $(a,b)$, exactly one of the following holds:
  1. (a)
    The expansion $\mathcal{U}_{a,b}$ is a reduct of a definitional expansion of Presburger arithmetic (possibly with countably many parameters), and in particular is NIP, dp-minimal, and does not interpret $(\mathbb{N},+,\times)$.
  2. (b)
    The expansion $\mathcal{U}_{a,b}$ interprets $(\mathbb{N},+,\times)$, and in particular has the independence property and is model-theoretically wild.
Moreover, case (b) does not occur: every Ulam expansion falls into case (a).
This conjecture asserts that the purely additive nature of the Ulam construction is incompatible with the encoding of multiplication, regardless of whether combinatorial rigidity holds. A proof would likely require new techniques combining additive combinatorics with definability theory.
Remark 7.27 (Relation to Zilber’s trichotomy).
Conjecture 7.26 is reminiscent of trichotomy and dichotomy principles in geometric model theory. In Zilber’s trichotomy for strongly minimal sets [Zilber1993], a strongly minimal set is either “trivial,” “group-like,” or “field-like,” with the latter corresponding to interpretability of algebraically closed fields. Our conjecture posits an analogous dichotomy for Ulam expansions: either the expansion remains “additively tame” or it becomes “arithmetically wild.” The content of the conjecture is that the second alternative never occurs for Ulam predicates.
7.4.6 Effectivity and computability considerations
We conclude this subsection with remarks on the effective content of the results under partial rigidity. The following is essentially a repackaging of the gap-regularity analysis in a more computational form.
Proposition 7.28 (Effective bounds under gap regularity).
Suppose $U(a,b)$ satisfies gap regularity (R2) with computable witnesses, i.e. there is an algorithm that, given $(a,b)$, outputs a threshold $N$ and period $p$ such that $g_{k+p}=g_{k}$ for all $k\geq N$. Then:
  1. (i)
    The density $d(U(a,b))$ is computable (as a rational number) from the witnesses.
  2. (ii)
    The theory $\mathrm{Th}(\mathcal{U}_{a,b})$ is decidable.
  3. (iii)
    Membership in any definable subset of $\mathbb{N}^{k}$ is decidable, uniformly in the defining formula.
Proof.
(i) From $N$ and $p$, compute $G=\sum_{j=0}^{p-1}g_{N+j}$ by simulating the Ulam construction up to stage $N+p$. Then $d(U(a,b))=p/G$ by Proposition 7.9(i).
(ii) By Proposition 7.9(ii), the predicate $\mathrm{U}_{a,b}$ is Presburger-definable with parameters determined by the witnesses. A formula in the language of $\mathcal{U}_{a,b}$ can therefore be translated into an equivalent Presburger formula (with the parameter values substituted). Presburger arithmetic is decidable (e.g. via Cooper’s algorithm or automata-theoretic methods), so the translated formula can be decided.
(iii) Given a formula $\varphi(\bar{x})$ in the language of $\mathcal{U}_{a,b}$ and a tuple $\bar{n}\in\mathbb{N}^{k}$, we can decide $\mathcal{U}_{a,b}\models\varphi(\bar{n})$ by the procedure in (ii). ∎
Remark 7.29.
The hypothesis of “computable witnesses” in Proposition 7.28 is stronger than mere truth of (R2). It is conceivable that $U(a,b)$ is gap-regular but that no algorithm can find the threshold $N$ and period $p$. In that case, the density and theory would still be determined (by absoluteness), but not necessarily effectively computable. The rigidity programme of Hinman–Kuca–Schlesinger–Sheydvasser [HinmanKucaSchlesingerSheydvasser2019JNT] aims in part to provide such computable witnesses for specific families.
8 Further directions and open problems
The results of this paper place a range of natural statements about Ulam sequences $U(a,b)$ at low levels of the arithmetical hierarchy, and use this to derive absoluteness and basic model-theoretic consequences for the expansions $(\mathbb{N},+,\mathrm{U}_{a,b})$. We end by listing some directions in which this picture could be refined or extended.
Refining complexity bounds
For rigidity, regularity and density, we have shown that the relevant statements are arithmetical and in fact lie at low levels of the hierarchy: strong rigidity and regularity are $\Sigma^{0}_{2}$ for fixed $(a,b)$, while family rigidity and density statements are $\Pi^{0}_{3}$ in the parameters. Several natural questions remain.
  • For fixed coprime $(a,b)$, are the $\Sigma^{0}_{2}$ and $\Pi^{0}_{3}$ classifications for rigidity, regularity, and density statements optimal in the sense of arithmetical reducibility? For example, does there exist a pair $(a,b)$ such that the corresponding rigidity sentence is $\Sigma^{0}_{2}$-complete (or the density sentence is $\Pi^{0}_{3}$-complete) under many-one or Turing reductions?
  • Can one show that certain weaker forms of rigidity, or certain regularity properties, are in fact equivalent to simpler $\Sigma^{0}_{1}$ or $\Pi^{0}_{1}$ sentences, perhaps after appropriate reformulation?
  • Conversely, are there natural Ulam-like constructions for which combinatorial conjectures give rise to statements of genuinely higher arithmetical complexity?
A more systematic proof-theoretic analysis along the lines of reverse mathematics (cf. Proposition 6.4) would also be of interest, for instance by calibrating global assertions such as $\mathrm{FRig}$ or uniform regularity principles between $\mathsf{RCA}_{0}$ and stronger systems.
Other Ulam-like constructions
The framework developed here applies to any additive construction whose membership relation is recursive and whose conjectured large-scale structure can be encoded by finite patterns with linear dependence on parameters. Natural candidates for similar analysis include:
  • Higher-dimensional Ulam sets in $\mathbb{Z}^{n}$, as studied by Kravitz–Steinerberger [KravitzSteinerberger2018] and Hinman–Kuca–Schlesinger–Sheydvasser [HinmanKucaSchlesingerSheydvasser2019Involve]. The arithmetisation methods extend naturally to recursively presented ordered abelian groups.
  • Variants in which the “unique representation” condition is replaced by bounded multiplicity (e.g. at most $k$ representations) or by weighted counting of representations of different types.
  • Other additive sequences such as sum-free sets, Sidon sets, and more general $B_{h}$-sequences, where membership is again defined via uniqueness or bounded multiplicity of representations as sums of previous terms. Many of these admit conjectural or proven interval-with-mask descriptions, and it would be natural to ask whether their rigidity and density statements enjoy analogous arithmetical and absoluteness properties.
  • Polynomial Ulam sequences in the sense of Sheydvasser [Sheydvasser2021LinearPolynomials], and more general algebraic deformations where the parameters $(a,b)$ are replaced by integer polynomials or higher-dimensional algebraic data.
For such systems, one can ask to what extent the sort of arithmetisation and complexity analysis developed here still applies, and whether similar absoluteness phenomena persist.
Model-theoretic classification
The expansion $(\mathbb{N},+,\mathrm{U}_{a,b})$ provides a concrete test case for the model theory of expansions of Presburger arithmetic by unary predicates. Under strong rigidity, we showed that $\mathcal{U}_{a,b}$ is a definitional expansion of Presburger arithmetic and hence is NIP and dp-minimal (Proposition 7.11). Several directions suggest themselves.
  • Can one prove NIP or other tameness properties of $\mathcal{U}_{a,b}$ under substantially weaker combinatorial hypotheses than full strong rigidity, for example assuming only regularity of the gap sequence together with mild growth conditions?
    Conjecture 8.1 (Tameness of all Ulam expansions).
    For every coprime $(a,b)$, the theory $\mathrm{Th}(\mathcal{U}_{a,b})$ is NIP and does not interpret $(\mathbb{N},+,\times)$. In particular, all expansions $(\mathbb{N},+,\mathrm{U}_{a,b})$ are model-theoretically tame in the sense of classification theory.
  • At the family level, can one obtain uniform bounds on VC-dimension or dp-rank for classes of sets definable in the multi-sorted structure obtained by varying $(a,b)$ and retaining a uniform predicate $\mathrm{U}(a,b,m)$?
More broadly, it would be interesting to relate the behaviour of Ulam sequences to existing general results on expansions of Presburger arithmetic by sparse sets or sets of prescribed asymptotic density.
Remark 8.2 (Automatic structures).
Expansions of $(\mathbb{N},+)$ whose definable sets are semilinear are closely related to automatic structures, where elements and relations are recognised by finite automata. Under strong rigidity, the predicate $\mathrm{U}_{a,b}$ is Presburger-definable, so $(\mathbb{N},+,\mathrm{U}_{a,b})$ sits very near this automatic realm. It would be interesting to make this connection precise, for example by asking whether suitable presentations of $\mathcal{U}_{a,b}$ are automatic in the sense of Khoussainov–Nerode, and whether the pattern decompositions arising from Ulam rigidity can be recovered by purely automata-theoretic means.
Broader logical context
Finally, the arithmetical nature of the Ulam problems suggests possible connections with other areas of logic. For example:
  • Since all the statements considered here are arithmetical, their truth values are invariant under set forcing and do not depend on the ambient set-theoretic universe. Thus any genuine independence phenomena would have to be of a proof-theoretic nature (for example, independence from a weak theory of arithmetic such as $\mathsf{RCA}_{0}$ or $\mathsf{PA}$), which are reflected in the behaviour of these statements in nonstandard models of arithmetic. This is conceptually distinct from the nonstandard methods used in [HinmanKucaSchlesingerSheydvasser2019JNT], where nonstandard extensions of $\mathbb{N}$ are employed as a tool inside ZFC to study the eventual behaviour of standard Ulam sequences.
  • One could also ask for categorical or topos-theoretic formulations of Ulam rigidity, for instance by interpreting the Ulam construction in suitable arithmetic or realizability toposes and analysing the extent to which the rigidity phenomena are preserved.
  • From the model-theoretic point of view, one might hope for analogues of the Marker–Steinhorn-type dichotomies [MarkerSteinhorn1994DefinableTypes] for expansions of Presburger arithmetic by sparse additive sets such as Ulam sequences.
We leave these questions for future work. Even at the purely arithmetical level, there remain many open combinatorial problems about Ulam sequences; the results of this paper show that, whatever their ultimate resolution, the logical complexity of the corresponding statements is tightly controlled.